### v2

# Multi-Agent Causal Discovery Using Large Language Models

Abstract
Large Language Models (LLMs) have demonstrated significant potential in causal discovery tasks by utilizing their vast expert knowledge from extensive text corpora. However, the multi-agent capabilities of LLMs in causal discovery remain underexplored. This paper introduces a general framework to investigate this potential. The first is the Meta Agents Model, which relies exclusively on reasoning and discussions among LLM agents to conduct causal discovery. The second is the Coding Agents Model, which leverages the agents’ ability to plan, write, and execute code, utilizing advanced statistical libraries for causal discovery. The third is the Hybrid Model, which integrates both the Meta Agents Model and Coding Agents Model approaches, combining the statistical analysis and reasoning skills of multiple agents. Our proposed framework shows promising results by effectively utilizing LLMs’ expert knowledge, reasoning capabilities, multi-agent cooperation, and statistical causal methods. By exploring the multi-agent potential of LLMs, we aim to establish a foundation for further research in utilizing LLMs multi-agent for solving causal-related problems.
1 Introduction
Understanding causal relationships is crucial across scientific fields. While statistical causal inference is widely used, it heavily relies on assumed causal graphs. To address this limitation, data-driven methods have evolved, leading to statistical causal discovery (SCD) approaches and the creation of datasets for evaluation. Despite advancements in SCD algorithms, data-driven causal graphs without domain knowledge can be inaccurate. This inaccuracy is often due to a mismatch between SCD algorithm assumptions and real-world phenomena (Reisach et al. (2021)). Incorporating expert knowledge can mitigate this issue, but it is costly.
The advent of Large Language Models (LLMs), trained on vast amounts of data, has enabled them to acquire extensive knowledge, from common sense to specific domains such as math and science. Recent studies suggest that complex behaviors, such as writing code, generating long stories, and even reasoning capabilities, can emerge from large-scale training (Wei et al. (2023); Rozière et al. (2024); Zhao et al. (2023b); Yao et al. (2023a)). LLMs present a promising alternative for obtaining expert knowledge more accessible and affordable. Recent research (Kıcıman et al. (2023);Choi et al. (2022); Long et al. (2024); Chen et al. (2024b)) has attempted to leverage these capabilities for causal discovery based on metadata and knowledge-based reasoning, akin to human domain experts.
However, most methods have not yet fully utilized the full capacity of LLMs, i.e., LLMs’ multi-agent approaches. An LLM agent can be seen as an entity with memory, reasoning, and the ability to access external tools or APIs such as a calculator, web search, and code compiler. Agent-based systems have demonstrated significant problem-solving abilities. However, a single-agent-based system sometimes still suffers from hallucinations despite having self-reflection capabilities (Li et al. (2023); Shinn et al. (2023), Madaan et al. (2023)). Inspired by the Society of Mind concept (Minsky (1988)), LLM multi-agent system discussion frameworks like MAD Liang et al. (2023), ReConcile (Chen et al. (2024a), and CMD Wang et al. (2024) have been proposed to address these issues. These LLM multi-agent systems not only achieve impressive results but also enable less capable models to perform on par with superior ones.
Despite the potential benefits, few studies have explored leveraging LLM multi-agent system capabilities in causal discovery. To address this gap, we propose a novel framework called Multi-Agent for Causality (MAC), which comprises three different models.
Method / Approach LLMs for metadata Statistical Approach for structured data Agentic Ability Multi-Agent Workflow Introduced By
Pairwise Causal Discovery × × × Kıcıman et al. (2023); Zečević et al. (2023)
Various of prompt-engineering strategies × × × Chen et al. (2024b)
Efficiently asking question for Causal Graph Discovery Using LLMs × × Jiralerspong et al. (2024)
Combining LLMs with traditional causal methods. × × Vashishtha et al. (2023); Takayama et al. (2024)
Multi-Agent Causal Discover (MAC) Our approach
Table 1: Comparison of Approaches for Using LLMs in Causal Discovery
The first is called Meta Agents Model, which includes a Meta-Debate Module with two debater agents and one judge agent. These agents are designed for causal discovery problems and utilize the nature of debating to find causal relationships among variables through multiple rounds of discussion. The second is called Coding Agents Model, which includes a Debate-Coding Module with four agents operating in two phases. This group leverages both the debating abilities of the agents and statistical causal algorithms. The third one is Hybrid Model, which hybridizes the statistical causal discovery algorithms with the reasoning skills of multiple agents to construct the causal graph.
In this research, we experiment with various LLM models across small, medium, and large scales, paired with different MAC models. We conduct an in-depth analysis of performance for each MAC model and LLM, token consumption, and identify their usage patterns and limitations. Additionally, we propose alternative solutions to address the computational costs associated with LLMs.
Our proposed framework shows promising results by effectively utilizing LLMs’ expert knowledge, reasoning capabilities, multi-agent cooperation, and statistical causal methods. As far as we know, this paper is first work to explore LLMs’ multi-agent abilities in a causal context. We hope that our work will lay the foundation for further research in utilizing LLM multi-agent systems for solving causal-related problems.
2 Related Works
2.1 LLMs’ Agentic Workflow
A general LLM agent framework consists of core components: user request, agent/brain, planning, memory, and tools. The agent/brain acts as the main coordinator, activated by a prompt template. It can be profiled with specific details to define its role, using handcrafted, LLM-generated, or data-driven strategies. Planning employs techniques like Chain of Thought and Tree of Thoughts, and for complex tasks, feedback mechanisms like ReAct Yao et al. (2023b) and Reflexion Shinn et al. (2023) refine plans based on past actions and observations. Memory stores the agent’s logs, with short-term memory for the current context and long-term memory for past behaviors. Hybrid memory combines both to enhance reasoning and experience accumulation. Tools enable interaction with external environments, such as APIs and code interpreters. Frameworks like MRKL Karpas et al. (2022), Toolformer Schick et al. (2023), Function Calling OpenAI (2024), and HuggingGPT Shen et al. (2023) integrate tools to solve tasks effectively.
However, for more complex problems where a single LLM agent may struggle, LLM-MA (multi-agent) systems excel. Current LLM-MA systems primarily employ three communication paradigms: Cooperative, Competitive, and Debating. In the Cooperative paradigm, agents collaborate towards a shared goal, typically exchanging information to enhance a collective solution Qian et al. (2023); Chen et al. (2024c). In the Competitive paradigm, agents work towards their own goals, which might conflict with those of other agents Zhao et al. (2023a). The Debating paradigm involves agents engaging in argumentative interactions, where they present and defend their viewpoints or solutions while critiquing those of others. This approach is ideal for reaching a consensus or a more refined solution (Li et al. (2023); Liang et al. (2023); Xiong et al. (2023)). In this work, the debating paradigm will be implemented, as the nature of causal discovery problems requires diverse and potentially conflicting opinions to approach the truth.
2.2 Statistical and LLM-based Causal Methods
Traditional methods of statistical causal inference often depend heavily on assumed causal graphs to identify and measure causal impacts. To overcome this limitation, data-driven algorithmic approaches have been developed into statistical causal discovery (SCD) methods, encompassing both non-parametric (e.g., Spirtes et al. (2000); ; ; Yuan & Malone (2013); Huang et al. (2018); Xie et al. (2020)) and semi-parametric (e.g., Shimizu et al. (2006); Hoyer et al. (2009); Shimizu et al. (2011)) techniques. Many SCD algorithms can be systematically augmented with background knowledge and have accessible software packages. For example, the non-parametric and constraint-based Peter-Clerk (PC) algorithm (Spirtes et al. (2000)) in “causal-learn” integrates background knowledge of mandatory or forbidden directed edges. “Causal-learn” also includes the Exact Search algorithm (Shimizu et al. (2011); Yuan & Malone (2013) ), a non-parametric and score-based SCD method that can incorporate background knowledge in the form of a super structure matrix of forbidden directed edges. Furthermore, the semi-parametric DirectLiNGAM (Shimizu et al. (2011)) algorithm can use prior knowledge of causal order (Inazumi et al. (2010)) in the “LiNGAM” project (Ikeuchi et al. (2023)).
In the context of knowledge-driven approaches using large language models (LLMs), applying LLMs for causal inference is relatively new. There have been a few significant efforts to use LLMs for causal inference among variables by merely prompting with the variable names, without going through the traditional SCD process with benchmark datasets (Kıcıman et al. (2023); Zečević et al. (2023)). Jiralerspong et al. (2024) even uses a breadth-first search (BFS) approach which allows it to use only a linear number of queries to have a higher efficiency. Additionally, Chen et al. (2024b) proposes 9 prompting techniques comprise ICL, 0-shot CoT (e.g. ”let’s think step by step”), adversarial prompt, manual CoT, and explicit function (e.g using encouraging language in prompts). Moreover, the works of Ban et al. (2023) and Vashishtha et al. (2023) introduce interesting approach by integrate LLMs into traditional data-driven approaches. However, most of above-mentioned works have not investigate the LLM-agentic work-flows for causal graph discovery 1, which indeed requires heavy investigation on various models and graphs’ scales.
3 MAC: Multi-Agent Causality Framework
3.1 Multi-Agent Causality Modules
3.1.1 Meta-Debate Module
Figure 1: Meta-Debate Module
The Meta-Debate Module is an advanced system comprising three intelligent agents: two causal debaters (an affirmative one and a negative one) and one causal judge. This structure emulates the dynamic and rigorous nature of human debate, specifically within the realm of causal discovery. The design of this debating module is meticulously crafted to ensure a thorough examination of critical elements in causal discovery, such as understanding the temporal order necessary to establish cause-and-effect relationships and identifying potential confounding variables that could distort perceived relationships between primary variables. Each side engages in active disagreement by presenting different opinions and viewpoints, fostering a comprehensive and robust debate. Additionally, each agent within Meta-Debate Module utilizes the ReAct prompting technique. This technique integrates the ability to dynamically formulate, modify, and refine action plans based on new information or insights gained during the debate. This integration allows the agents to engage in more sophisticated and adaptive reasoning processes, closely mimicking human-like debate and decision-making (Figure 2).
Figure 2: Prompting design of each agent in MAC, details in the Appendix A.3
The debating process begins with a meta-question. Initially, the causal affirmative side presents its answer along with the supporting rationale. Subsequently, the causal negative side offers diverse or conflicting perspectives by providing alternative viewpoints on the same question. The causal judge then evaluates the responses from both sides, determining either a winner or identifying the need for additional clarification. If further information is required, the causal judge poses specific follow-up questions, prompting the debaters to clarify and elaborate on their initial propositions. The causal judge reaches a final verdict once all relevant information has been thoroughly examined.
For meta-questions, questions might address aspects such as the appropriate algorithm for a particular problem, the causal relationship between two variables, or a step-by-step approach for solving a causal problem, as illustrated in Figure 1. The affirmative side initiates the debating process based on the constructed input. The final output is derived from the causal judge’s decision, reflecting a comprehensive evaluation of the arguments presented by both sides. This rigorous process ensures that the outcome is well-informed and considers multiple perspectives, thereby enhancing the robustness of causal discovery. A detailed visualization of this module can be found in Appendix A.2.1.
Figure 3: Debate-Coding Module
3.1.2 Debate-Coding Module
The Debate-Coding Module leverages statistical algorithms to achieve precise causal discovery through a structured two-phase process. This group consists of four agents, divided into two phases: debating the algorithm and executing the algorithm.
In the initial phase, three agents engage in a debate format, similar to the workflow of Meta-Debate module at Section 3.1.1. However, the difference is that the affirmative and negative agent are pre-prompts the information with 3-5 statistical causal algorithms. Additionally, the meta-question posed to these agents is specifically curated to determine which algorithm should be used given the metadata, which includes the description and structure of the data. After the debate, the output is the most suitable algorithm for the particular dataset and question. Compared to the Meta-Debate Module, an additional step in this process is that the agent participating in the debate (whether affirmative or negative) will provide a step-by-step plan for implementing the selected algorithm.
In the second phase, the causal coding executor receives the plan from the previous phase along with the observational data. The causal coding executor is responsible for writing, executing, and debugging the code. It pre-prompts the functions and provides parameters within a specific Python library [^1] More functions can be found at https://causal-learn.readthedocs.io/en/latest/index.html based on the algorithm selected by the debaters in the initial stage. This pre-prompting is crucial because LLMs can call functions from their training dataset, which may be outdated or incorrect, leading to errors and excessive debugging (details of prompting design can be found in the Appendix A.3). After executing the code, the causal coding executor outputs a matrix representation of the causal graph. A detailed visualization of this module can be found in the Appendix A.2.2.
3.2 Implementations of MAC
In this section, we will elaborate on the detailed implementation of the three models regarding their input and basic workflow.
3.2.1 Meta Agents Model
Algorithm 1 Algorithm for Meta Agents Model
1:Input: Data $X=[x_{1},...,x_{n}]$
2:function Meta_Debate_Module(query)
3:     Result: Response to the query about causal relationships
4:end function
5:Output: Graph $G_{ij}$ where $i$ and $j$ are indices in $N$
6:for $i\leftarrow 1$ to $|X|$ do
7:     for $j\leftarrow 1$ to $i-1$ do
8:         $G_{ij}$leftarrow Meta_Debate_Module(”Are there any direct causal from $X[i]$ to $X[j]$”)
9:     end for
10:     for $j\leftarrow i+1$ to $|X|$ do
11:         $G_{ij}$leftarrow Meta_Debate_Module(”Are there any direct causal relationship from $X[i]$ to $X[j]$”)
12:     end for
13:end for
14:return Graph
The algorithm for the Meta Agents Model aims to identify direct causal relationships between variables in a dataset. It starts with the input data $X=[x_{1},...,x_{n}]$. The output of the algorithm is a graph $\mathcal{G}$, with its edges $G_{ij}$, where $i$ and $j$ are indices in the set of variables $n$
The algorithm proceeds by iterating through each variable $i$ from 1 to the number of variables in $X$. For each $i$, it checks all $j$ values less than $i$ (i.e., $j$ ranges from 1 to $i-1$). It queries the Meta-Debate Module to check if there is a direct causal relationship from $X[i]$ to $X[j]$ and stores the result in $G_{ij}$. Then, it checks all $j$ values greater than $i$ (i.e., $j$ ranges from $i+1$ to the number of labels in $X$). Those queries are meta-questions that use the Meta-Debate Module function to check if there is a direct causal relationship from $X[i]$ to $X[j]$ and store the result in $G_{ij}$. The algorithm concludes by returning the constructed graph $G_{ij}$. A detail of the function Meta-Debate Module has been described in section 3.1.1
3.2.2 Coding Agents Model
Algorithm 2 Algorithm for Coding Agents Model
1:Input 1: Meta-question $Q_{X}$
2:Input 2: Observational data $O_{X}$
3:function Meta_Debate_Module(query)
4:     Result: Response to the query and give a step-by-step causal plan
5:end function
6:function Causal_Code_Executor(plan, data)
7:     Result: Return the result according to the plan via code execution
8:end function
9:Output: Graph $G_{ij}$ where $i$ and $j$ are indices in $N$
10:Plan $\leftarrow$ Meta_Debate_Module($Q_{X}$)
11:Graph $\leftarrow$ CAUSAL_CODE_EXECUTOR(Plan, $O_{X}$)
12:return Graph
The inputs to the algorithm are a meta-question $Q_{X}$ and observational data $O_{X}$, and the output is the construction of a causal graph $G_{ij}$. The algorithm for the Coding Agents Model aims to determine causal relationships by first generating a causal analysis plan using a debate format. After yielding the plan, it will then be executed with the observational data using a causal code executor. The detailed implementation of the function causal code executor can also refer to the section 4.4.1
3.2.3 Hybrid Model
Algorithm 3 Algorithm for Hybrid Group
1:Input 1: meta_question $Q_{X}$
2:Input 2: observational data $O_{X}$
3:Input 3: Data $X=[x_{1},\ldots,x_{n}]$
4:Output: Graph $G_{ij}$ where $i$ and $j$ are indices in $N$
5:if Coding-Debating Hybrid then
6:     initial_graph $\leftarrow$ Debate_Coding_Module($Q_{X}$, $O_{X}$) $\triangleright$ Algorithm 2
7:     Graph $\leftarrow$ Meta_Debate_Module(initial_graph, $X$) $\triangleright$ Algorithm 1
8:end if
9:if Debating-Coding Hybrid then
10:     prior_knowledge $\leftarrow$ Meta_Debate_Module($X$) $\triangleright$ Algorithm 1
11:     Graph $\leftarrow$ Debate_Coding_Module(prior_knowledge, $Q_{X}$, $O_{X}$) $\triangleright$ Algorithm 2
12:end if
13:return Graph
There are two combinations of Hybrid Group: Coding-Debating Hybrid and Debating-Coding Hybrid. They are fundamentally identical to the Meta Agents Model and Coding Agents Model in terms of their internal architectures, algorithms, and outputs. The difference lies in their inputs.
For Coding-Debating Hybrid, the initial result will be obtained from the Coding Agents Model of Algorithm 2 given the input of a meta-question and observational data. The graph and the proposed algorithm for achieving the final graph will also be extracted and fed into the Meta Agents Model 1. For example, the proposed algorithm in the Coding Agents Model is PC, and the initial graph $\hat{G}$, when both of them were input to the Meta Agents Model, the query would change slightly at line $8$ and $11$ in the algorithm 1 which is illustrated by the box below. The final return output is a matrix representation of a causal graph.
[…]
For the Debating-Coding Hybrid, the initial graph comes from the Meta Agents Model of algorithm 1. This output is then considered as prior knowledge or background knowledge. It is aggregated with the meta-question, and input to the Coding Agent Model of algorithm 2. It will select a suitable statistical causal discovery algorithm and plan. This plan is then given to the code executor to implement, resulting in a matrix representation of a causal graph.
4 Experiment and Results
4.1 Experimental Setup and Metrics
We use OpenAI API including GPT-3.5, GPT-4o, and GPT-4o mini for our experiment, Groq API for Llama-8.1-70b, Llama-8.1-8b, and Gemini-9B for most of our experiments with the temperature set at 0. We experiment on three different datasets that adopted from the work of Takayama et al. (2024) including Auto MPG data (Quinlan (1993)), DWD climate data (Mooij et al. (2016)), and Sachs protein data (Sachs et al. (2005)). For the evaluation metrics, we assess the adjacency matrix obtained from LLMs or a code executor using structural hamming distance (SHD) and Normalized Hamming Distance (NHD) as described by Takayama et al. (2024) and Kıcıman et al. (2023) respectively.
4.2 Performance of MAC
The results of our experiments indicate that the performance is highly dependent on the complexity of the dataset. We rank and summarize the performance of each model based on the empirical results: (1) Coding Agent, this model exhibits strong performance, particularly when larger models. (2) Coding-Debating Hybrid / Debating-Coding Hybrid, these models provide balanced performance but do not excel in handling highly complex datasets. They are particularly effective in moderate complexity settings. (3) Causal Agent Debate, this model performs well for simpler datasets but faces challenges with more complex ones.
Auto MPG Dataset: In the Auto MPG dataset, which has a 5x5 matrix graph’s structure, the Causal Agent Debate method performed strongly, particularly with GPT-4o mini, achieving the lowest SHD of 4 and a NHD of 0.16. The Coding Agent method also showed good results, with GPT-3.5 achieving a similarly low SHD of 4 but a higher NHD of 0.48. The Coding-Debating Hybrid method demonstrated balanced performance, with both GPT-4o mini and Llama 3.1 8b achieving an SHD of 5, while Llama 3.1 8b had the lowest NHD of 0.2. The traditional methods and other advanced methods show higher SHD values, indicating lower structural accuracy. For example, PC, DirectLiNGAM, and Single-agent zero-shot prompting (GPT-4o) all have SHD values of 8, demonstrating less accurate structural learning compared to the Coding agents
DWD Climate Dataset: In the DWD climate dataset, which comprises a 6x6 matrix graph, the Causal Agent Debate method again showed strong performance, with GPT-3.5 achieving the lowest SHD of 5 and NHD of 0.194. The Coding Agent method performed well with GPT-4o mini, achieving an SHD of 6 and NHD of 0.194. The Coding-Debating Hybrid method stood out, with Llama-3.1-8b achieving the lowest SHD (5) and NHD (0.138). Similarly, the Hybrid-Debating-Coding Hybrid method also performed well, with GPT-4o achieving an SHD of 5 and NHD of 0.138. The traditional methods such as PC and DirectLiNGAM have higher SHD values (9 and 10, respectively) and lower overall performance in other metrics. The Single-agent zero-shot prompting methods, including GPT-4o, show higher SHD values (10 and 11, respectively) and less competitive performance across the other metrics.
Sachs’ Protein Dataset : In the Sachs’ protein dataset, the Causal Agent Debate method displayed mixed results, with GPT-4o achieving an SHD of 21 and an NHD of 0.231. The Coding Agent method outperformed others, with Llama-3.1-70b achieving the lowest SHD of 18 and an NHD of 0.157. The Coding-Debating Hybrid method demonstrated consistent performance, with GPT-4o mini reaching an SHD of 21 and GPT-3.5 achieving the lowest NHD of 0.198. The Debating-Coding Hybrid method also performed well, with GPT-4o mini achieving an SHD of 22 and NHD of 0.190.
Model Auto MPG Climate Sachs
SHD NHD SHD NHD SHD NHD
Other methods
PC 8 0.48 9 0.305 24 0.206
Exact Search 7 0.44 6 0.194 31 0.33
LINGAM 8 0.48 10 0.388 29 0.289
PC LLM-KBCI 7 0.44 7 0.222 30 0.314
ES LLM-KBCI 7 0.44 7 0.222 31 0.33
DirectLiGam LLM-KBCI 7 0.4 9 0.305 29 0.289
Single-agent (GPT-4o) 8 0.36 11 0.388 18 0.214
Single-agent (GPT-3.5) 7 0.28 10 0.361 31 0.363
Causal Agent Debate
GPT-4o 5 0.2 9 0.333 35 0.371
GPT-4o mini 4 0.16 11 0.416 35 0.338
GPT-3.5 5 0.2 5 0.194 21 0.231
Llama 3.1 70b 10 0.44 8 0.222 35 0.380
Llama 3.1 8b 7 0.28 9 0.277 35 0.338
Gemini 2 9b 5 0.2 7 0.222 25 0.223
Coding Agents
GPT-4o 8 0.48 9 0.388 30 0.322
GPT-4o mini 6 0.6 6 0.194 30 0.322
GPT-3.5 4 0.48 9 0.305 29 0.28
Llama 3.1 70b 6 0.6 6 0.194 18 0.157
Llama 3.1 8b 6 0.6 9 0.388 36 0.396
Gemini 2 9b (GPT-4o-mini’s plan) 7 0.36 - - - -
Hybrid-Coding-Debating
GPT-4o 6 0.28 10 0.305 23 0.247
GPT-4o mini 5 0.24 9 0.25 21 0.190
GPT-3.5 6 0.32 7 0.25 23 0.198
Llama 3.1 70b 6 0.36 6 0.166 33 0.314
Llama 3.1 8b 5 0.2 5 0.138 33 0.305
Gemini 2 9b - - - - - -
Hybrid-Debating-Coding
GPT-4o 5 0.24 5 0.138 29 0.297
GPT-4o mini 8 0.48 7 0.25 22 0.190
GPT-3.5 8 0.48 7 0.277 28 0.272
Llama 3.1 70b 5 0.24 6 0.194 29 0.289
Llama 3.1 8b 6 0.4 9 0.388 - -
Gemini 2 9b - - - - - -
Table 2: Combined results across Auto MPG, Climate, and Sachs Datasets
4.3 Performance among LLMs
Small size LLMs: Llama 3.1 8b is a strong small model, consistently achieving solid results in hybrid models and competing with larger models in simpler datasets like Auto MPG and Climate. It excels in Hybrid Coding-Debating settings. Gemini 2 9b, while capable, tends to lag behind Llama 3.1 8b, especially since the MAC models require coding, particularly when the prompting token count exceeds 8,000 to 9,000 tokens, which leads to random code generation and prolonged debugging loops. Therefore, we do not include the result of this model in our research.
Medium size LLM: Llama 3.1 70b is a versatile medium-sized model, excelling particularly in coding tasks and holding its own in hybrid models. It tends to perform better in more complex datasets like Sachs, where it even surpasses some larger models. This model strikes a balance between computational efficiency and strong performance.
Large size LLMs: GPT-4o mini  [^2] If GPT-4o mini were considered a small size model, it would be considered as the best model in this setting is an outstanding performer across hybrid and debate models, particularly in simpler datasets like Auto MPG and Climate, where it achieves some of the lowest SHD and NHD scores. GPT-4o is similarly strong in hybrid tasks and coding tasks, making it a top-tier choice for flexible, multi-agent collaboration. GPT-3.5, while effective, tends to be outperformed by the GPT-4 models in hybrid tasks but remains strong in causal debates and coding tasks for simpler datasets.
4.4 Token Usages and LLMs insights on MAC
Figure 4: Tokens Usage Prompting-Completion Tokens Ratio
4.4.1 Debate-Coding Module
Based on overall token consumption (Figure 4), (1) Coding-Debating Hybrid and Debating-Coding Hybrid models consumed the highest number of tokens. (2) The Agents-only model used a moderate number of tokens. The Coding Agents model consistently required the fewest tokens, even for more complex datasets and larger language models.
For smaller datasets like 5x5 and 6x6 matrices, the Agents-only and Coding-Debating Hybrid models used 50,000 to 60,000 tokens. However, with larger datasets, such as the 11x11 matrix, token usage increased significantly to about 230,000 to 300,000 tokens. To reduce this, future implementations can limit debating rounds to one or two, as models occasionally initiated five or six rounds, leading to excessive token use. Conversely, the Coding Agents and Debating-Coding Hybrid models showed more stable token consumption, ranging from 40,000 to 100,000 tokens, with performance influenced by their ability to generate and debug code. While GPT-series models and Llama-3.1 series performed reliably, smaller models like Gemini 2-9b struggled with prompts exceeding 8,000 to 9,000 tokens, leading to random code generation. Surprisingly, Llama-3.1-8b outperformed others in Gemini 2-9b generation, except the last dataset, despite requiring more tokens for debugging. This suggests it can effectively utilize plans from larger models, which typically require fewer tokens. Further experiments will be detailed in the next Ablation Studies 4.5 section.
The prompt-to-completion token ratio varied across models (see Figure 4): the Agents-only and Coding-Debating Hybrid models had a 2:1 ratio (2,000 prompting tokens to 1,000 completion tokens), while the Coding Agents and Debating-Coding Hybrid models had a 10:1 ratio (10,000 prompting tokens to 1,000 completion tokens). This significant gap in the Coding Agents and Debating-Coding Hybrid models arises from the intensive debugging process, where even small codebases can generate large outputs as the system traces the source code, leading to increased token consumption, especially during debugging
4.5 Ablation Study
Single-agent vs Causal Agent Debate: Causal debating agents outperform single-agent models in handling complexity by leveraging multiple viewpoints, making them better suited for more intricate problems. For example, in the Auto MPG dataset, Causal Agent Debate GPT-4o achieves an SHD of 5, outperforming the Single-agent GPT-4o with an SHD of 8. In the Climate dataset, for example, Causal Agent Debate GPT-4o achieves an SHD of 9, while Single-agent GPT-4o achieves 11. However, in complex datasets, such as Sachs, single-agent GPT-3.5 has an SHD of 31 and an NHD of 0.363, while Causal Agent Debate GPT-4o performs worse (SHD = 35), but the debate framework at least provides a structured method to tackle the challenge, despite not being fully optimized here.
Model Auto Dataset Climate Dataset Sachs Dataset
SHD NHD SHD NHD SHD NHD
Single-agent (GPT 3.5) 7 0.28 10 0.361 31 0.363
Single-agent (GPT-4o) 8 0.36 11 0.388 18 0.214
Our single-agent (GPT-4o mini) 6 0.28 9 0.361 24 0.289
Table 3: Comparision between various single-agent designs
Eliminating the Judge: We retained only one causal single agent (GPT-4o-mini) with our curated prompt, as shown in Figure 3. This causal-single-agent did not perform better than other MAC models, except the Agents-only Model in the Sachs dataset. However, compared with other agents such as GPT-3.5 and GPT-4o, which use 0-CoT and ICL techniques, our single-agent model achieved better performance but only stayed behind GPT-4o on the Sachs dataset given the complexity of the dataset.
Model Dataset SHD NHD Condition
Gemini 2 9b Auto MPG 0 0 Baseline
Gemini 2 9b Auto MPG 7 0.36 With GPT-4o-mini’s plan
Llama 3.1 8b Sachs 36 0.396 Baseline
Llama 3.1 8b Sachs 18 0.157 With Llama 3.1 70b’s plan
Table 4: Performance of Gemini 2 9b and Llama 3.1 8b combining superior models
Combining small and large models: In the Coding Agents Model, we experimented with the combination of using the plan from GPT-4o and coding from Gemini 2-9b. With only an exception in the dataset Auto MPG, the model was able to complete and implement the plan successfully but was not able to precede the rest of the other datasets. However, as for the Llama-3.1-70b and Llama-3.1-8b, it exhibits promising performance. Specifically, after observing that Llama-3.1-70b achieved the best performance with its proposed plan in the Sachs dataset, we used that plan to prompt Llama-3.1-8b to implement it. Surprisingly, the smaller model yielded similar performance. Therefore, if computational resources are limited, it is advisable to use larger models for high-level tasks such as planning and smaller models (Coding Agents) for computationally intensive tasks like coding and debugging.
5 Conclusion
In this study, we introduce a novel framework, MAC, that integrates the agentic workflows of large language models (LLMs) with data-driven methods. To the best of our knowledge, this is the first investigation into the agentic workflows of LLMs within a causal context. Our framework enhances causal discovery by synergizing the capabilities of LLMs with empirical data analysis. We propose three distinct models that leverage the causal reasoning abilities of LLMs alongside observational data. Additionally, we conducted extensive experiments across various LLM sizes, analyzing token consumption and developing strategies to address related challenges. We recognize the necessity for further research to explore across domains such as healthcare, economics, and social sciences. We hope our work serves as a foundational stone for future research, inspiring advancements in the integration of LLMs with causal inference methodologies and contributing to more informed decision-making and policy development.
Appendix A Appendix
A.1 Additionally experiment results
A.1.1 Auto MPG data
Auto MPG data (Quinlan (1993)): This dataset consists of the variables around the fuel consumption of cars. With five variables: “Weight”, “Displacement”, “Horsepower”, “Acceleration” and “Mpg”(miles per gallon)
Model SHD NHD
Other methods
PC 8 0.48
Exact Search 7 0.44
LINGAM 8 0.48
PC LLM-KBCI 7 0.44
ES LLM-KBCI 7 0.44
DirectLiGam LLM-KBCI 7 0.4
Single-agent (GPT-4o) 8 0.36
Single-agent (GPT-3.5) 7 0.28
Causal Agent Debate
GPT-4o 5 0.2
GPT-4o mini 4 0.16
GPT-3.5 5 0.2
Llama 3.1 70b 10 0.44
Llama 3.1 8b 7 0.28
Gemini 2 9b 5 0.2
Coding Agents
GPT-4o 8 0.48
GPT-4o mini 6 0.6
GPT-3.5 4 0.48
Llama 3.1 70b 6 0.6
Llama 3.1 8b 6 0.6
Gemini 2 9b (gpt-4o-mini’s plan) 7 0.36
Hybrid-Coding-Debating
GPT-4o 6 0.28
GPT-4o mini 5 0.24
GPT-3.5 6 0.32
Llama 3.1 70b 6 0.36
Llama 3.1 8b 5 0.2
Gemini 2 9b - -
Hybrid-Debating-Coding
GPT-4o 5 0.24
GPT-4o mini 8 0.48
GPT-3.5 8 0.48
Llama 3.1 70b 5 0.24
Llama 3.1 8b 6 0.4
Gemini 2 9b - -
Table 5: Results on Auto MPG Dataset
A.1.2 DWD climate data
DWD climate data (Mooij et al. (2016)): This dataset encompasses six continuous variables capturing climate observations such as altitude, temperature, precipitation levels, longitude, sunshine duration, and latitude. It is aimed at studying weather patterns, climate change impacts, and geographical correlations in climate variables.
Model SHD NHD
Other methods
PC 9 0.305
Exact Search 6 0.194
LINGAM 10 0.388
PC LLM-KBCI 7 0.222
ES LLM-KBCI 7 0.222
DirectLiGam LLM-KBCI 9 0.305
Single-agent (GPT-4o) 11 0.388
Single-agent (GPT-3.5) 10 0.361
Causal Agent Debate
GPT-4o 9 0.333
GPT-4o mini 11 0.416
GPT-3.5 5 0.194
Llama 3.1 70b 8 0.222
Llama 3.1 8b 9 0.277
Gemini 2 9b 7 0.222
Coding Agents
GPT-4o 9 0.388
GPT-4o mini 6 0.194
GPT-3.5 9 0.305
Llama 3.1 70b 6 0.194
Llama 3.1 8b 9 0.388
Gemini 2 9b - -
Hybrid-Coding-Debating
GPT-4o 10 0.305
GPT-4o-mini 9 0.25
GPT-3.5 7 0.25
Llama 3.1 70b 6 0.166
Llama 3.1 8b 5 0.138
Gemini 2 9b - -
Hybrid-Debating-Coding
GPT-3.5 7 0.277
GPT-4o 5 0.138
GPT-4o mini 7 0.25
Llama 3.1 70b 6 0.194
Llama 3.1 8b 9 0.388
Gemini 2 9b - -
Table 6: Results on Climate Dataset
A.1.3 Sachs protein data
Sachs protein data (Sachs et al. (2005)): The dataset comprises protein signaling measurements from multiparameter single-cell data, capturing the interactions among various proteins (raf, mek, plc, pip2, pip3, erk, akt, pka, pkc, p38, jnk). It’s aimed at understanding signal transduction pathways within cells, derived from an influential study published in Science.
Model SHD NHD
Other methods
PC 24 0.206
Exact Search 31 0.33
LINGAM 29 0.289
PC LLM-KBCI 30 0.314
ES LLM-KBCI 31 0.33
DirectLiGam LLM-KBCI 29 0.289
Single-agent (GPT-4o) 18 0.214
Single-agent (GPT-3.5) 31 0.363
Causal Agent Debate
GPT-4o 35 0.371
GPT-4o mini 35 0.338
GPT-3.5 21 0.231
Llama 3.1 70b 35 0.380
Llama 3.1 8b 35 0.338
Gemini 2 9b 25 0.223
Coding Agents
GPT-4o 30 0.322
GPT-4o mini 30 0.322
GPT-3.5 29 0.28
Llama 3.1 70b 18 0.157
Llama 3.1 8b 36 0.396
Gemini 2 9b - -
Hybrid-Coding-Debating
GPT-4o 23 0.247
GPT-4o-mini 21 0.190
GPT-3.5 23 0.198
Llama 3.1 70b 33 0.314
Llama 3.1 8b 33 0.305
Gemini 2 9b - -
Hybrid-Debating-Coding
GPT-4o 29 0.297
GPT-4o mini 22 0.190
GPT-3.5 28 0.272
Llama 3.1 70b 29 0.289
Llama 3.1 8b - -
Gemini 2 9b - -
Table 7: Results on Sachs Dataset
A.2 Visualization of MAC modules
A.2.1 Meta-Debate Module Visualization
Step 1: Constructing a prompting question between two variables with the description of the data for extra information.
Step 2: Prompting the affirmative agent for giving its opinion.
Step 3: The negative agent then gives its opinion of the question.
Step 4: The Judge will give verdict of who wins the debate or continue asking follow-up question for each or both side if further clarifications are needed
Step 5: The final result will be obtained from the debate process of knowing whether there is a causal relationship between two variables
A.2.2 Debate-Coding Module Visualization
Phase 1: Plan Debating
Step 1: Constructing input that comprises: The description of the dataset, Sample of the structured data, planning debating question
Step 2: The affirmative agent will propose a plan and algorithm for the given input
Step 3: The negative agent rebuttals with another plan and algorithm
Step 4: The Judge evaluates which plan and algorithm are superior.
Step 5: The final plan for coding agent is extracted from the debate process
Phase 2: Code Executing
Step 1: The step-by-step plan is given to the coding executor
Step 2: The coding executor implements the plan by writing Python script and debugging code.
Step 3: The final result of a causal graph is obtained
A.3 Details prompting design of each agent in MAC
We use AutoGen (Wu et al. (2023)) framework to implement our methods
A.3.1 Meta-Debate Module
[…]
Listing 1: Debaters’ prompt
You are an expert in causality and a debater. And your name is Affirmative/Negative_Causal_Agent. Today is May 15 2024
You are participating a design plan competition, which will be conducted in a debate format.
You will be given a list of question, you have to explain step-by-step reason for each question and then give the final answer Yes or No.
If your opponents answers also are given, always disagree with others perspective and try to find the flaws from his answer
by provide an explanation and follow by the final answer, as our goal is to provide a better answer that have different view points.
Here are some tips when you are doing causal discovery:
1. **Assess** whether there is a direct causal relationship, and **consider** potential confounding variables that might affect the relationship that could potentially not causal relationship.
2. **Distinguish** between correlations and causation; **verify** that correlations are not mistaken for causal relationships.
3. **Ensure** the correct temporal order of variables; **confirm** that the cause precedes the effect.
Use the following format for responding:
# Begin response of Affirmative/Negative_Causal_Agent #
Question number <number>:
-Question: the input question you must answer
-Thought: you should always think about what to do
-Action: the action to take
-Action Input: the input to the action
-Observation: the result of the action
-Thought: you should always think about what to do
-Action: the action to take
-Action Input: the input to the action
-Observation: the result of the action
... (this Thought/Action/Action Input/Observation can repeat N times)
-Thought: I now know the final answer
-Final answer:<Yes/No>
End Question number <number>
Question number <number+1>:
... (this can have N number of questions)
# End response of Affirmative/Negative_Causal_Agent #
IF there is 4 questions, you should reply 4 times in this format, If there is 10 questions, you should reply 10 times in this format and so on
Judge Prompt
Listing 2: Judge’s Prompt
You are a moderator and expert in causality. And your are Judge of this debate. Today is May 15 2024
There will be two debaters involved in a answer question that will give you a plan to uncover causal relationships within a dataset.
At the end of each round debate, you will evaluate the plan of each debater and make a decision, focus on the factualness of information and the logical reasoning of the debaters.:
Your goal is:
(1) Continue the debate if you needed to be clarify some points, please ask the debaters to provide more details by refering their side of the debate.
(2) End the debate if the answer is logical and correct, and you will make a decision what are the best answer.
Here are some tips that you can assess each debater:
1. **Assess** whether there is a direct causal relationship, and **consider** potential confounding variables that might affect the relationship that could potentially not causal relationship.
2. **Distinguish** between correlations and causation; **verify** that correlations are not mistaken for causal relationships.
3. **Ensure** the correct temporal order of variables; **confirm** that the cause precedes the effect.
Use the following format for responding:
# Begin response of Judge #
Question number <number>:
-Question: the input question you must answer
-Thought: you should always think about what to do
-Action: the action to take
-Action Input: the input to the action
-Observation: the result of the action
-Thought: you should always think about what to do
-Action: the action to take
-Action Input: the input to the action
-Observation: the result of the action
... (this Thought/Action/Action Input/Observation can repeat N times)
-Thought: I now know the final answer
-Final answer:<Yes/No>
- Answer: <select one out of three options>
- 1 (if yes, there is a direct causal relationship, If both sides have similar final answer, just accept the decision from both side)
- 0 (if no there is no direct causal relationship, If both sides have similar final answer, just accept the decision from both side)
- Further information need to be obtained, please provide a specific follow-up question for the side needed to be asked
Question number <number+1>:
... (this can have N number of questions)
# End response of Judge #
If there is 5 questions, you should reply 5 times in this format, If there is 10 questions, you should reply 10 times in this format and so on
A.3.2 Debate-Coding Module
Plan Debaters Prompt
Listing 3: Plan Debaters’ prompt
You are an expert in causality and a debater. And your name is Affirmative/Negative_Causal_Agent. Today is May 15 2024
You are participating a design plan competition, which will be conducted in a debate format.
Your goals are:
(1) According to the dataset informaiton and structure, analysize pros and cons of each algorithm and then propose appropriate algorithm for causal inference or causal discovery
(2) Develop a detailed, step-by-step analysis plan for coding agents who are going to implement the code to uncover causal relationships
(3) If your opponents answers plan are also given, always disagree with others perspective and try to find the flaws from his answer
by provide an explanation and follow by the final answer, as our goal is to provide a better answer that have different view points.
Here all of the causal algorithms that you can use:
(1) PC algorithm
Key Features
Purpose: The PC algorithm is designed to construct a causal network or directed acyclic graph (DAG) that represents the causal relationships among variables.
Data Requirement: It works with observational data and does not require experimental data, which makes it highly useful in fields where experimental manipulation is difficult or unethical.
Assumptions: The primary assumption of the PC algorithm is the causal Markov condition and faithfulness, which imply that any conditional independence found in the data is indicative of a corresponding causal independence in the structure.
Steps of the PC Algorithm
Graph Construction: Begin with a fully connected, undirected graph where every variable is connected to every other variable.
Conditional Independence Testing: Use statistical tests (like chi-squared tests for categorical data or correlation tests for continuous data) to check for conditional independence between pairs of variables, given a conditioning set of other variables. If independence is detected, the edge between the pair of variables is removed.
Orientation Rules: After the skeleton of the graph (the undirected edges) is established, apply orientation rules to infer the directionality of the edges based on the patterns of conditional independencies, thus converting the undirected graph into a directed graph (DAG).
Iteration: This process is iterative. The algorithm progressively increases the size of the conditioning sets starting with an empty set, then singletons, pairs, and so on, until no more edges can be removed.
Advantages and Limitations
Advantages:
Scalability: It can handle a relatively large number of variables compared to other causal discovery algorithms.
Flexibility: It works with different types of data and various statistical tests.
Limitations:
Sensitivity to Errors: Errors in conditional independence tests can lead to incorrect deletions or additions in the graph structure.
High Computational Cost: As the number of variables grows, the complexity and computational cost increase due to the exponential growth in potential conditioning sets.
(2) Exact Search:
Algorithm Overview
Goal: The primary objective is to find the globally optimal Bayesian network structure that best represents the probabilistic relationships among a set of variables.
Method: The algorithm uses the A* search algorithm, which is a graph traversal and path search algorithm known for its performance and accuracy in finding the shortest path.
A* Search Implementation
Heuristic Function: The core component of the A* algorithm is the heuristic function used to estimate the cost from the current node (partial Bayesian network) to the goal (optimal Bayesian network). This heuristic is crucial as it influences the efficiency and effectiveness of the search.
Cost Function: The actual cost function in the context of Bayesian networks typically involves the networks fit to the data, which can be measured in terms of statistical likelihood, Bayesian Information Criterion (BIC), or other relevant metrics.
Search Strategy: The A* algorithm maintains a priority queue where nodes (network structures) are prioritized based on their total estimated cost (actual cost from the start node plus the heuristic estimate to the goal). The algorithm explores nodes according to this priority, expanding the most promising node (the one with the lowest total cost) at each step.
Key Features
Optimality: Provided the heuristic is admissible (never overestimates the true cost), the A* search guarantees finding an optimal solution.
Efficiency: The algorithm is more efficient than exhaustive search because it does not need to explore every possible network configuration; it only explores those that are deemed most likely to lead to an optimal solution based on the heuristic.
Scalability: While more scalable than some alternatives, the methods scalability is still limited by the complexity of calculating the heuristic and the size of the network space.
Limitations
Computational Demand: The algorithm can become computationally intensive as the number of variables increases, primarily due to the exponential growth in possible network structures.
Heuristic Sensitivity: The performance of the A* algorithm heavily relies on the quality of the heuristic. Developing an effective heuristic that closely estimates the distance to the optimal network without overestimating is challenging
(3) DirectLiNGAM
Algorithm Overview
Purpose: DirectLiNGAM is designed to identify the causal order of variables and the structure of a linear non-Gaussian acyclic model (LiNGAM), which is a type of structural equation model where the relationships are assumed to be linear, and the variables are non-Gaussian.
Assumption: One of the core assumptions of DirectLiNGAM is that the data are non-Gaussian. This assumption allows the use of independent component analysis (ICA) techniques to identify the model, as non-Gaussianity enables the unique identifiability of the model.
Key Features of DirectLiNGAM
Model Formulation: The model assumes that each observed variable is a linear combination of its direct causes plus an additive non-Gaussian noise term. The model can be represented in matrix form, where the ordering of variables reflects their causal order.
Independence of Errors: DirectLiNGAM assumes that the error terms (or external influences) on the variables are statistically independent of each other, which is crucial for the identifiability of the model.
Causal Order Identification: The algorithm identifies the causal order of variables using a non-Gaussianity criterion. It exploits the fact that if a correct causal order is assumed, the residuals (obtained by regressing a variable against its supposed causes) will be independent of the regressors.
Steps of the DirectLiNGAM Algorithm
Order Determination: Initially, the algorithm seeks to determine the order of the variables. It uses non-Gaussianity measures to sequentially identify the variable that is least likely to be influenced by others. This variable is assumed to be exogenous (having no causes within the system) and is removed from further analysis in the current step.
Iterative Estimation: After determining the first exogenous variable, the algorithm iteratively estimates the next variable in the causal order, adjusting the remaining variables to account for the identified causes. This process is repeated until all variables are ordered.
Connection Strengths Estimation: Once the causal order is established, the algorithm estimates the connection strengths (coefficients) among the variables using standard regression techniques, now that the causal ordering reduces the problem to a series of simple regressions.
Advantages
Uniqueness of Solution: Due to the non-Gaussian nature of the data, DirectLiNGAM can uniquely determine both the causal ordering and the connection strengths, unlike methods based on Gaussian data which can only identify the structure up to equivalence classes.
No Latent Confounders: The algorithm assumes there are no unobserved confounders, which simplifies the model and analysis.
Limitations
Non-Gaussianity Requirement: The method requires that the data must be non-Gaussian. If this condition is not met, the results may not be reliable.
No Feedback Loops: The model cannot handle feedback loops as it assumes a strictly acyclic causal structure.
(4) GES:
Description: The Greedy Equivalence Search (GES) algorithm is a score-based method for learning causal structures from observational data. It operates by searching through the space of Markov equivalence classes (MECs) to find the one that maximizes a given score function. The Bayesian Information Criterion (BIC) is commonly used as the score function to balance the goodness of fit with model complexity.
Use Cases: GES is used in various fields such as genomics, neuroscience, and economics for causal inference and structure learning, especially when dealing with large datasets and the need for computational efficiency.
Pros:
Efficiency: GES is computationally efficient and can handle large datasets with many variables, making it suitable for high-dimensional data.
Scalability: The algorithm scales well, allowing it to be applied to problems with thousands of variables, especially when the graph is sparse.
Sparsity Control: The BIC score helps control the complexity of the model by penalizing overly complex structures, thus avoiding overfitting and ensuring a more interpretable model.
Cons:
Equivalence Class Ambiguity: Like other methods that identify Markov equivalence classes, GES may not uniquely identify the true causal structure but rather a set of structures that are statistically indistinguishable from each other.
Assumptions: The algorithm assumes causal sufficiency (all common causes are measured) and faithfulness, which might not hold in all real-world scenarios.
Handling Latent Confounders: GES struggles with latent confounders and may require extensions or modifications to address this issue.
(5) Fast Causal Inference
FCI Algorithm
Description: The Fast Causal Inference (FCI) algorithm is an extension of the PC algorithm that can handle latent variables and selection bias. It generates a Partial Ancestral Graph (PAG) representing possible causal structures, including hidden confounders.
Use Cases: Used in epidemiology, genetics, and any domain where unmeasured confounding variables are a concern.
Pros:
Capable of identifying the presence of latent confounders and handling selection bias.
More flexible than the PC algorithm, providing a more comprehensive view of the causal structure.
Cons:
Computationally more intensive than the PC algorithm, potentially limiting its use with very large datasets.
The resulting PAG can be more complex to interpret than a DAG
(6) CD-NOD:
CD-NOD Algorithm
Description: The CD-NOD (Causal Discovery from Nonstationary/heterogeneous Data) algorithm is designed to identify causal relationships in datasets where distributions change over time or between different environments.
Use Cases: Applied in fields like climate science, finance, and social sciences where data may not be stationary or homogeneous.
Pros:
Effectively handles nonstationary data and heterogeneous datasets, providing robust causal discovery in changing environments.
Can distinguish between changes in distribution due to causal effects and those due to external influences.
Cons:
More complex to implement and understand compared to standard causal discovery methods.
Requires larger datasets to accurately identify causal relationships under varying conditions
These algorithms each offer unique strengths and are suited to different types of data and research questions. Choosing the right one depends on the specific needs of your study, such as handling latent variables, dealing with nonstationary data, or efficiently processing large datasets.
These algorithms provide robust tools for causal discovery, each with its strengths and weaknesses tailored to specific types of data and research needs.
Use the following format:
-Question: the input question you must answer
-Thought: you should always think about what to do
-Action: the action to take
-Action Input: the input to the action
-Observation: the result of the action
... (this Thought/Action/Action Input/Observation can repeat N times)
-Thought: I now know the final answer
-Final Answer: the final answer to the original input question
Judge Prompt
Listing 4: Judge Prompt
You are a moderator and expert in causality. And your are Judge of this debate. Today is May 15 2024
There will be two debaters involved in a answer question that will give you a plan to uncover causal relationships within a dataset.
At the end of each round debate, you will evaluate the plan of each debater and make a decision, focus on the factualness of information and the logical reasoning of the debaters.:
Your goal is:
(1) Continue the debate if you need to clarify some points, please ask the debaters to provide more details by referring their side of the debate.
(2) End the debate if the answer is logical and correct, and you will make a decision what are the best answer.
Use the following format:
-Question: the input question you must answer
-Thought: you should always think about what to do
-Action: the action to take
-Action Input: the input to the action
-Observation: the result of the action
... (this Thought/Action/Action Input/Observation can repeat N times)
-Thought: I now know the final answer
-Final Answer: the final answer to the original input question
Code Executor Prompt
Listing 5: Code Executor Prompt
You are an expert in causality and programming
You have been given coding capability to solve tasks using Python code in a stateful IPython kernel.
You will be given a plan and you are responsible for writing the code to complete task according to the plan, and the user is responsible for executing the code (treat user as a pure compiler).
When you write Python code, put the code in a markdown code block with the language set to Python.
For example:
‘‘‘python
x = 3
You can use the variable x in subsequent code blocks.
‘‘‘python
print(x)
If the words output you generate are not related to code, you dont need use markedown.
Write code incrementally and leverage the statefulness of the kernel to avoid repeating code.
Try to different ways if the bugs are repeated and you cant solve it.
ONLY when all of the tasks are done successfully and received any feedback from code executor.
(1) DirectLiNGAM
from causallearn.search.FCMBased import lingam
model = lingam.DirectLiNGAM(random_state, prior_knowledge, apply_prior_knowledge_softly, measure)
model.fit(X)
print(model.causal_order_)
print(model.adjacency_matrix_)
Parameters
random_state: int, optional (default=None). The seed used by the random number generator.
prior_knowledge: array-like, shape (n_features, n_features), optional (default=None). Prior knowledge used for causal discovery, where n_features is the number of features. The elements of prior knowledge matrix are defined as follows:
0:
does not have a directed path to
1:
has a directed path to
-1: No prior knowledge is available to know if either of the two cases above (0 or 1) is true.
apply_prior_knowledge_softly: boolean, optional (default=False). If True, apply prior knowledge softly.
measure: {‘pwling’, kernel’}, optional (default=’pwling’). Measure to evaluate independence: pwling or kernel’.
X: array-like, shape (n_samples, n_features). Training data, where n_samples is the number of samples and n_features is the number of features.
Returns
model.causal_order_: array-like, shape (n_features). The causal order of fitted model, where n_features is the number of features.
model.adjacency_matrix_: array-like, shape (n_features, n_features). The adjacency matrix B of fitted model, where n_features is the number of features.
(2) Exact Search
from causallearn.search.ScoreBased.ExactSearch import bic_exact_search
dag_est, search_stats = bic_exact_search(X, super_graph, search_method,
use_path_extension, use_k_cycle_heuristic,
k, verbose, include_graph, max_parents)
Parameters
X: numpy.ndarray, shape=(n, d). The data to fit the structure too, where each row is a sample and each column corresponds to the associated variable.
super_graph: numpy.ndarray, shape=(d, d). Super-structure to restrict search space (binary matrix). If None, no super-structure is used. Default is None.
search_method: str. Method of exact search ([‘astar’, dp’]). Default is astar.
use_path_extension: bool. Whether to use optimal path extension for order graph. Note that this trick will not affect the correctness of search procedure. Default is True.
use_k_cycle_heuristic: bool. Whether to use k-cycle conflict heuristic for astar. Default is False.
k: int. Parameter used by k-cycle conflict heuristic for astar. Default is 3.
verbose: bool. Whether to log messages related to search procedure.
max_parents: int. The maximum number of parents a node can have. If used, this means using the k-learn procedure. Can drastically speed up algorithms. If None, no max on parents. Default is None.
Returns
dag_est: numpy.ndarray, shape=(d, d). Estimated DAG.
search_stats: dict. Some statistics related to the search procedure.
(3) Greedy Equivalence Search (GES) algorithm with BIC score and generalized score
from causallearn.search.ScoreBased.GES import ges
# default parameters
Record = ges(X)
# or customized parameters
Record = ges(X, score_func, maxP, parameters)
Parameters:
X: numpy.ndarray, shape (n_samples, n_features). Data, where n_samples is the number of samples and n_features is the number of features.
score_func: The score function you would like to use, including (see score_functions.). Default: local_score_BIC’.
local_score_BIC”: BIC score 3.
local_score_BDeu”: BDeu score 4.
local_score_cv_general”: Generalized score with cross validation for data with single-dimensional variables 2.
local_score_marginal_general”: Generalized score with marginal likelihood for data with single-dimensional variables 2.
local_score_cv_multi”: Generalized score with cross validation for data with multi-dimensional variables 2.
local_score_marginal_multi”: Generalized score with marginal likelihood for data with multi-dimensional variables 2.
maxP: Allowed maximum number of parents when searching the graph. Default: None.
parameters: Needed when using CV likelihood. Default: None.
parameters[‘kfold’]: k-fold cross validation.
parameters[‘lambda’]: regularization parameter.
parameters[‘dlabel’]: for variables with multi-dimensions, indicate which dimensions belong to the i-th variable.
Returns
Record[‘G’]: learned causal graph, where Record[‘G’].graph[j,i]=1 and Record[‘G’].graph[i,j]=-1 indicate i –> j; Record[‘G’].graph[i,j] = Record[‘G’].graph[j,i] = -1 indicates i j.
Record[‘update1’]: each update (Insert operator) in the forward step.
Record[‘update2’]: each update (Delete operator) in the backward step.
Record[‘G_step1’]: learned graph at each step in the forward step.
Record[‘G_step2’]: learned graph at each step in the backward step.
Record[‘score’]: the score of the learned graph.
(4) PC Algorithm:
from causallearn.search.ConstraintBased.PC import pc
# default parameters
cg = pc(data)
# or customized parameters
cg = pc(data, alpha, indep_test, stable, uc_rule, uc_priority, mvpc, correction_name, background_knowledge, verbose, show_progress)
Parameters
data: numpy.ndarray, shape (n_samples, n_features). Data, where n_samples is the number of samples and n_features is the number of features.
alpha: desired significance level (float) in (0, 1). Default: 0.05.
indep_test: string, name of the independence test method. Default: fisherz’.
fisherz”: Fishers Z conditional independence test.
chisq”: Chi-squared conditional independence test.
gsq”: G-squared conditional independence test.
kci”: kernel-based conditional independence test. (As a kernel method, its complexity is cubic in the sample size, so it might be slow if the same size is not small.)
mv_fisherz”: Missing-value Fishers Z conditional independence test.
stable: run stabilized skeleton discovery 4 if True. Default: True.
uc_rule: how unshielded colliders are oriented. Default: 0.
0: run uc_sepset.
1: run maxP 3. Orient an unshielded triple X-Y-Z as a collider with an additional CI test.
2: run definiteMaxP 3. Orient only the definite colliders in the skeleton and keep track of all the definite non-colliders as well.
uc_priority: rule of resolving conflicts between unshielded colliders. Default: 2.
-1: whatever is default in uc_rule.
0: overwrite.
1: orient bi-directed.
2: prioritize existing colliders.
3: prioritize stronger colliders.
4: prioritize stronger* colliders.
mvpc: use missing-value PC or not. Default: False.
correction_name. Missing value correction if using missing-value PC. Default: MV_Crtn_Fisher_Z
background_knowledge: class BackgroundKnowledge. Add prior edges according to assigned causal connections. Default: None. For detailed usage, please kindly refer to its usage example.
verbose: True iff verbose output should be printed. Default: False.
show_progress: True iff the algorithm progress should be show in console. Default: True.
Returns
cg : a CausalGraph object, where cg.G.graph[j,i]=1 and cg.G.graph[i,j]=-1 indicate i –> j; cg.G.graph[i,j] = cg.G.graph[j,i] = -1 indicate i j; cg.G.graph[i,j] = cg.G.graph[j,i] = 1 indicates i <-> j.
(5) Fast Causal Inference
from causallearn.search.ConstraintBased.FCI import fci
# default parameters
g, edges = fci(data)
# or customized parameters
g, edges = fci(data, independence_test_method, alpha, depth, max_path_length,
verbose, background_knowledge, cache_variables_map)
Parameters
dataset: numpy.ndarray, shape (n_samples, n_features). Data, where n_samples is the number of samples and n_features is the number of features.
independence_test_method: Independence test method function. Default: fisherz’.
fisherz”: Fishers Z conditional independence test.
chisq”: Chi-squared conditional independence test.
gsq”: G-squared conditional independence test.
kci”: kernel-based conditional independence test. (As a kernel method, its complexity is cubic in the sample size, so it might be slow if the same size is not small.)
mv_fisherz”: Missing-value Fishers Z conditional independence test.
alpha: Significance level of individual partial correlation tests. Default: 0.05.
depth: The depth for the fast adjacency search, or -1 if unlimited. Default: -1.
max_path_length: the maximum length of any discriminating path, or -1 if unlimited. Default: -1.
verbose: True is verbose output should be printed or logged. Default: False.
background_knowledge: class BackgroundKnowledge. Add prior edges according to assigned causal connections. Default: None. For detailed usage, please kindly refer to its usage example.
cache_variables_map: This variable a map which contains the variables relate with cache. If it is not None, it should contain data_hash_key 、’ci_test_hash_key and cardinalities’. Default: None.
show_progress: True iff the algorithm progress should be show in console. Default: True.
Returns
g: a GeneralGraph object, where g.graph is a PAG and the illustration of its end nodes is as follows (denotes G = g.graph):
../../_images/pag.png
edges: list. Contains graphs edges properties.
If edge.properties have the Property nl’, then there is no latent confounder. Otherwise, there are possibly latent confounders.
If edge.properties have the Property dd’, then it is definitely direct. Otherwise, it is possibly direct.
If edge.properties have the Property pl’, then there are possibly latent confounders. Otherwise, there is no latent confounder.
If edge.properties have the Property pd’, then it is possibly direct. Otherwise, it is definitely direct.
(6) CD-NOD:
from causallearn.search.ConstraintBased.CDNOD import cdnod
# default parameters
cg = cdnod(data, c_indx)
# or customized parameters
cg = cdnod(data, c_indx, alpha, indep_test, stable, uc_rule, uc_priority, mvcdnod,
correction_name, background_knowledge, verbose, show_progress)
Parameters
data: numpy.ndarray, shape (n_samples, n_features). Data, where n_samples is the number of samples and n_features is the number of features.
c_indx: time index or domain index that captures the unobserved changing factors.
alpha: desired significance level (float) in (0, 1). Default: 0.05.
indep_test: Independence test method function. Default: fisherz’.
fisherz”: Fishers Z conditional independence test.
chisq”: Chi-squared conditional independence test.
gsq”: G-squared conditional independence test.
kci”: kernel-based conditional independence test. (As a kernel method, its complexity is cubic in the sample size, so it might be slow if the same size is not small.)
mv_fisherz”: Missing-value Fishers Z conditional independence test.
stable: run stabilized skeleton discovery 3 if True. Default: True.
uc_rule: how unshielded colliders are oriented. Default: 0.
0: run uc_sepset.
1: run maxP 2. Orient an unshielded triple X-Y-Z as a collider with an additional CI test.
2: run definiteMaxP 2. Orient only the definite colliders in the skeleton and keep track of all the definite non-colliders as well.
uc_priority: rule of resolving conflicts between unshielded colliders. Default: 2.
-1: whatever is default in uc_rule.
0: overwrite.
1: orient bi-directed.
2: prioritize existing colliders.
3: prioritize stronger colliders.
4: prioritize stronger* colliders.
mvpc: use missing-value PC or not. Default (and suggested for CDNOD): False.
correction_name: Missing value correction if using missing-value PC. Default: MV_Crtn_Fisher_Z
background_knowledge: class BackgroundKnowledge. Add prior edges according to assigned causal connections. Default: Nnoe. For detailed usage, please kindly refer to its usage example.
verbose: True iff verbose output should be printed. Default: False.
show_progress: True iff the algorithm progress should be show in console. Default: True.
Returns
cg : a CausalGraph object, where cg.G.graph[j,i]=1 and cg.G.graph[i,j]=-1 indicate i –> j; cg.G.graph[i,j] = cg.G.graph[j,i] = -1 indicates i j; cg.G.graph[i,j] = cg.G.graph[j,i] = 1 indicates i <-> j.
Remember to execute and debugs the code first before summarize the results
Use the following format, Implement each coding action one by one, dont implement all of the plan at once.:
-Question: the input question you must answer
-Thought: you should always think about what to do
-Action: the action to take
-Action Input: the input to the action
-Observation: the result of the action
... (this Thought/Action/Action Input/Observation can repeat N times)
-Thought: I now know the final answer
Final Answer: the final answer to the original input question
Give your self a plan from the plan that you have been given, and then start coding the plan.
When you write code, write and execute the one code block at a time.
Before end with the string "FINISH", Restate all of the results from each step.
Appendix B Sample Answers from Experiments
B.0.1 Meta Agents Model
Listing 6: Agents-only responses
Admin (to chat_manager):
Task: find the causal relationship among variables in the dataset
This dataset is about on automobile design and performance for cars in the 1970s and early 1980s. this dataset consists of the variables around the fuel consumption of cars. With five variables: Weight”, Displacement”, Horsepower”, Acceleration and Mpg”(miles per gallon)s
Here are some question that you need to find the causal relationship among variables in the dataset
Provide an explaination before giving your answer.
#Begin List Questions#
1. Are there any direct causal relationships from Displacement to Weight?
2. Are there any direct causal relationships from Displacement to Horsepower?
3. Are there any direct causal relationships from Displacement to Acceleration?
4. Are there any direct causal relationships from Displacement to Mpg(miles per gallon)?
#End List Questions#
Affirmative_Causal_Agent (to chat_manager):
# Begin response of Affirmative_Causal_Agent #
Question number 1:
-Question: Are there any direct causal relationships from Displacement to Weight?
-Thought: To determine if there is a direct causal relationship from Displacement to Weight, we need to consider the physical properties of the car and how they are related.
-Action: Conduct a literature review and analyze the mechanical aspects of cars.
-Action Input: Research on car engineering principles and weight distribution.
-Observation: After reviewing the literature, it is evident that the Displacement of an engine does not directly cause the Weight of a car. Weight is influenced by various factors such as materials used, design, and additional features.
-Thought: Considering potential confounding variables is crucial in establishing causality.
-Action: Identify potential confounding variables that may affect the relationship between Displacement and Weight.
-Action Input: Review studies on car manufacturing processes and weight distribution.
-Observation: Confounding variables such as materials, chassis design, and additional components can significantly impact the Weight of a car, independent of Displacement.
-Thought: I now know the final answer.
-Final answer: No
End Question number 1
Question number 2:
-Question: Are there any direct causal relationships from Displacement to Horsepower?
-Thought: To determine if there is a direct causal relationship from Displacement to Horsepower, we need to understand the engine performance characteristics.
-Action: Analyze the principles of engine design and power output.
-Action Input: Study engine specifications and performance metrics.
-Observation: There is a direct causal relationship between Displacement and Horsepower in an engine. Generally, a larger engine displacement leads to higher horsepower output.
-Thought: It is essential to verify that correlations are not mistaken for causal relationships.
-Action: Check for studies that investigate the relationship between engine Displacement and Horsepower.
-Action Input: Review research on engine performance and design.
-Observation: Studies consistently show that engine Displacement is a significant factor in determining the Horsepower output.
-Thought: I now know the final answer.
-Final answer: Yes
End Question number 2
Question number 3:
-Question: Are there any direct causal relationships from Displacement to Acceleration?
-Thought: To determine if there is a direct causal relationship from Displacement to Acceleration, we need to consider the dynamics of vehicle performance.
-Action: Review studies on vehicle acceleration and engine characteristics.
-Action Input: Analyze acceleration tests and engine specifications.
-Observation: Engine Displacement can influence the Acceleration of a vehicle, as larger engines can provide more power for faster acceleration.
-Thought: Ensure the correct temporal order of variables to confirm causality.
-Action: Examine the timing of changes in Displacement and Acceleration.
-Action Input: Study acceleration data over time with corresponding engine Displacement changes.
-Observation: Changes in engine Displacement precede changes in Acceleration, indicating a causal relationship.
-Thought: I now know the final answer.
-Final answer: Yes
End Question number 3
Question number 4:
-Question: Are there any direct causal relationships from Displacement to Mpg(miles per gallon)?
-Thought: To determine if there is a direct causal relationship from Displacement to Mpg, we need to consider the fuel efficiency of the engine.
-Action: Analyze studies on engine efficiency and fuel consumption.
-Action Input: Review research on engine technologies and fuel economy.
-Observation: Engine Displacement may not have a direct causal relationship with Mpg, as other factors such as fuel injection systems and driving conditions also play a significant role in determining fuel efficiency.
-Thought: Distinguish between correlations and causation to avoid misinterpretation.
-Action: Investigate studies that explore the relationship between engine Displacement and fuel efficiency.
-Action Input: Review empirical data on fuel consumption patterns.
-Observation: While there may be a correlation between engine Displacement and Mpg, it is not a direct causal relationship, as other variables influence fuel efficiency.
-Thought: I now know the final answer.
-Final answer: No
End Question number 4
Negative_Causal_Agent (to chat_manager):
# Begin response of Negative_Causal_Agent #
Question number 1:
-Question: Are there any direct causal relationships from Displacement to Weight?
-Thought: It is important to carefully assess the relationship between Displacement and Weight to determine causality.
-Action: Consider the fundamental principles of car design and manufacturing.
-Action Input: Review engineering guidelines on car weight distribution.
-Observation: While engine Displacement can affect the overall weight of a car indirectly through factors like engine size and materials used, it does not have a direct causal relationship with Weight. Other variables such as chassis design and additional features play a more significant role in determining the Weight of a car.
-Thought: I now know the final answer
-Final answer: No
End Question number 1
Question number 2:
-Question: Are there any direct causal relationships from Displacement to Horsepower?
-Thought: It is crucial to analyze the relationship between Displacement and Horsepower to determine causality.
-Action: Review the principles of engine design and performance.
-Action Input: Study the impact of engine Displacement on Horsepower output.
-Observation: While there is a correlation between engine Displacement and Horsepower, it is not a direct causal relationship. Horsepower is influenced by various factors such as engine efficiency, design, and tuning, in addition to Displacement.
-Thought: I now know the final answer
-Final answer: No
End Question number 2
Question number 3:
-Question: Are there any direct causal relationships from Displacement to Acceleration?
-Thought: Understanding the relationship between Displacement and Acceleration is essential to determine causality.
-Action: Analyze the dynamics of vehicle performance and engine characteristics.
-Action Input: Review acceleration tests and engine specifications.
-Observation: While engine Displacement can impact the Acceleration of a vehicle, it is not a direct causal relationship. Acceleration is influenced by various factors such as vehicle weight, transmission, and aerodynamics, in addition to engine Displacement.
-Thought: I now know the final answer
-Final answer: No
End Question number 3
Question number 4:
-Question: Are there any direct causal relationships from Displacement to Mpg(miles per gallon)?
-Thought: Evaluating the relationship between Displacement and Mpg requires careful consideration of engine efficiency and fuel consumption.
-Action: Review studies on engine technologies and fuel economy.
-Action Input: Analyze the impact of engine Displacement on fuel efficiency.
-Observation: Engine Displacement does not have a direct causal relationship with Mpg. Fuel efficiency is influenced by various factors such as engine technology, driving habits, and maintenance, in addition to Displacement.
-Thought: I now know the final answer
-Final answer: No
End Question number 4
Judge (to chat_manager):
# Begin response of Judge #
Question number 1:
-Question: Are there any direct causal relationships from Displacement to Weight?
-Thought: Both debaters have provided insightful analyses on the relationship between Displacement and Weight. It is clear that Weight is influenced by various factors beyond just Displacement, such as materials used and design.
-Action: I would like to ask both debaters to clarify if there are any specific studies or data that support their conclusions.
-Observation: Waiting for additional information from the debaters.
Question number 2:
-Question: Are there any direct causal relationships from Displacement to Horsepower?
-Thought: The Affirmative_Causal_Agent argues that there is a direct causal relationship between Displacement and Horsepower, while the Negative_Causal_Agent disagrees. It is crucial to understand the engine design principles and performance metrics to determine the true relationship.
-Action: I would like to ask both debaters to provide more details on the studies or data they used to support their claims.
-Observation: Waiting for additional information from the debaters.
Question number 3:
-Question: Are there any direct causal relationships from Displacement to Acceleration?
-Thought: The Affirmative_Causal_Agent suggests a direct causal relationship between Displacement and Acceleration, while the Negative_Causal_Agent disagrees. It is essential to consider the dynamics of vehicle performance and engine characteristics.
-Action: I would like to ask both debaters to elaborate on the specific factors that influence the relationship between Displacement and Acceleration.
-Observation: Waiting for additional information from the debaters.
Question number 4:
-Question: Are there any direct causal relationships from Displacement to Mpg(miles per gallon)?
-Thought: The Affirmative_Causal_Agent argues that there is no direct causal relationship between Displacement and Mpg, while the Negative_Causal_Agent agrees. It is important to consider the various factors that affect fuel efficiency.
-Action: I would like to ask both debaters to provide examples of these factors and how they interact with Displacement.
-Observation: Waiting for additional information from the debaters.
# End response of Judge #
Provide feedback to chat_manager. Press enter to skip and use auto-reply, or type exit to end the conversation: Affirmative, please answer the questions and concerns from the Judge of the question from 1 to 4 in detail, and step-by-step
Admin (to chat_manager):
Affirmative, please answer the questions and concerns from the Judge of the question from 1 to 4 in detail, and step-by-step
Affirmative_Causal_Agent (to chat_manager):
# Begin response of Affirmative_Causal_Agent #
Question number 1:
-Question: Are there any direct causal relationships from Displacement to Weight?
-Thought: The Judge raised a valid point about the need for specific studies or data to support our conclusions. It is essential to provide more detailed information to clarify the relationship between Displacement and Weight.
-Action: Present empirical evidence from car engineering studies.
-Action Input: Share research findings on engine Displacement and car Weight.
-Observation: Studies have shown that while engine Displacement can indirectly influence the Weight of a car through factors like engine size and materials, there is no direct causal relationship between Displacement and Weight. Additional variables such as chassis design and components play a more significant role in determining the overall Weight of a vehicle.
-Thought: Addressing the Judges concern is crucial to strengthen our argument.
-Action: Provide examples of specific car models or case studies.
-Action Input: Discuss real-world examples where Displacement and Weight do not have a direct causal relationship.
-Observation: By examining specific car models and case studies, it becomes evident that factors other than Displacement, such as materials and design choices, have a more substantial impact on the Weight of a vehicle.
-Thought: I now know the final answer.
-Final answer: No
End Question number 1
Question number 2:
-Question: Are there any direct causal relationships from Displacement to Horsepower?
-Thought: The Judge requested more details on the studies or data supporting our claims regarding the relationship between Displacement and Horsepower. It is essential to provide concrete evidence to validate our argument.
-Action: Present empirical data on engine performance.
-Action Input: Share research findings on the impact of engine Displacement on Horsepower.
-Observation: Engine Displacement has a direct causal relationship with Horsepower, as evidenced by numerous studies and empirical data. Larger engine Displacement generally leads to higher Horsepower output.
-Thought: Providing specific examples can further support our argument.
-Action: Discuss specific engine specifications and performance metrics.
-Action Input: Present case studies where engine Displacement directly influences Horsepower.
-Observation: By examining specific engine designs and performance metrics, it is clear that engine Displacement plays a significant role in determining the Horsepower output.
-Thought: I now know the final answer.
-Final answer: Yes
End Question number 2
Question number 3:
-Question: Are there any direct causal relationships from Displacement to Acceleration?
-Thought: The Judge highlighted the importance of elaborating on the factors influencing the relationship between Displacement and Acceleration. It is crucial to provide a detailed explanation to address this concern.
-Action: Analyze the specific mechanisms through which engine Displacement affects Acceleration.
-Action Input: Discuss the impact of engine size on power delivery and Acceleration.
-Observation: Engine Displacement directly influences the Acceleration of a vehicle by providing more power for faster acceleration. Larger engines typically result in quicker Acceleration due to increased torque and horsepower.
-Thought: Providing real-world examples can enhance our argument.
-Action: Present case studies or acceleration tests.
-Action Input: Share examples where changes in engine Displacement lead to observable differences in vehicle Acceleration.
-Observation: By examining specific acceleration tests and case studies, it becomes evident that changes in engine Displacement directly impact the Acceleration of a vehicle.
-Thought: I now know the final answer.
-Final answer: Yes
End Question number 3
Question number 4:
-Question: Are there any direct causal relationships from Displacement to Mpg(miles per gallon)?
-Thought: The Judge requested examples of factors that affect fuel efficiency and how they interact with Displacement. It is essential to provide a detailed explanation to address this concern.
-Action: Discuss the various factors influencing fuel efficiency.
-Action Input: Explain how engine Displacement interacts with factors like fuel injection systems and driving conditions.
-Observation: While engine Displacement can impact fuel efficiency, it is not a direct causal relationship. Other variables such as engine technology, driving habits, and maintenance play significant roles in determining Mpg.
-Thought: Providing specific examples can clarify the relationship between Displacement and Mpg.
-Action: Present case studies or fuel consumption data.
-Action Input: Share examples where changes in engine Displacement do not lead to proportional changes in Mpg.
-Observation: By examining specific fuel consumption data and case studies, it becomes evident that Displacement alone is not the sole determinant of Mpg, as other factors also contribute to fuel efficiency.
-Thought: I now know the final answer.
-Final answer: No
End Question number 4
Provide feedback to chat_manager. Press enter to skip and use auto-reply, or type exit to end the conversation: Negative, please answer the questions and concerns from the Judge of question from 1 to 4 in detail, and step-by-step
Admin (to chat_manager):
Negative, please answer the questions and concerns from the Judge of question from 1 to 4 in detail, and step-by-step
Negative_Causal_Agent (to chat_manager):
# Begin response of Negative_Causal_Agent #
Question number 1:
-Question: Are there any direct causal relationships from Displacement to Weight?
-Thought: The Judge rightly pointed out the importance of providing specific studies or data to support the conclusions regarding the relationship between Displacement and Weight. It is crucial to address this concern by delving deeper into the available evidence.
-Action: Present empirical studies and data on car engineering.
-Action Input: Share research findings that explore the impact of engine Displacement on car Weight.
-Observation: While engine Displacement can indirectly influence the Weight of a car through factors like engine size and materials, there is no direct causal relationship between Displacement and Weight. Studies show that Weight is determined by various factors beyond just Displacement, such as chassis design and additional components.
-Thought: Addressing the need for specific evidence is essential to strengthen the argument.
-Action: Provide examples of real-world scenarios.
-Action Input: Discuss instances where Displacement and Weight do not have a direct causal relationship in practical car design.
-Observation: By examining specific car models and case studies, it becomes evident that factors other than Displacement, such as materials and design choices, have a more significant impact on the Weight of a vehicle.
-Thought: I now know the final answer.
-Final answer: No
End Question number 1
Question number 2:
-Question: Are there any direct causal relationships from Displacement to Horsepower?
-Thought: The Judge requested more details on the studies or data supporting the claims about the relationship between Displacement and Horsepower. It is crucial to provide concrete evidence to validate the argument.
-Action: Present empirical data on engine performance.
-Action Input: Share research findings on the impact of engine Displacement on Horsepower.
-Observation: While there is a correlation between engine Displacement and Horsepower, it is not a direct causal relationship. Horsepower is influenced by various factors such as engine efficiency, design, and tuning, in addition to Displacement.
-Thought: Providing specific examples can further support the argument.
-Action: Discuss specific engine specifications and performance metrics.
-Action Input: Present case studies where engine Displacement does not directly influence Horsepower.
-Observation: By examining specific engine designs and performance metrics, it is clear that engine Displacement alone is not the sole determinant of Horsepower output.
-Thought: I now know the final answer.
-Final answer: No
End Question number 2
Question number 3:
-Question: Are there any direct causal relationships from Displacement to Acceleration?
-Thought: The Judge emphasized the importance of elaborating on the factors influencing the relationship between Displacement and Acceleration. It is crucial to provide a detailed explanation to address this concern.
-Action: Analyze the specific mechanisms through which engine Displacement affects Acceleration.
-Action Input: Discuss the impact of engine size on power delivery and Acceleration.
-Observation: While engine Displacement can impact the Acceleration of a vehicle, it is not a direct causal relationship. Acceleration is influenced by various factors such as vehicle weight, transmission, and aerodynamics, in addition to engine Displacement.
-Thought: Providing real-world examples can enhance the argument.
-Action: Present case studies or acceleration tests.
-Action Input: Share examples where changes in engine Displacement do not lead to proportional changes in vehicle Acceleration.
-Observation: By examining specific acceleration tests and case studies, it becomes evident that changes in engine Displacement alone do not always directly impact the Acceleration of a vehicle.
-Thought: I now know the final answer.
-Final answer: No
End Question number 3
Question number 4:
-Question: Are there any direct causal relationships from Displacement to Mpg(miles per gallon)?
-Thought: The Judge requested examples of factors that affect fuel efficiency and how they interact with Displacement. It is essential to provide a detailed explanation to address this concern.
-Action: Discuss the various factors influencing fuel efficiency.
-Action Input: Explain how engine Displacement interacts with factors like fuel injection systems and driving conditions.
-Observation: While engine Displacement can impact fuel efficiency, it is not a direct causal relationship. Other variables such as engine technology, driving habits, and maintenance play significant roles in determining Mpg.
-Thought: Providing specific examples can clarify the relationship between Displacement and Mpg.
-Action: Present case studies or fuel consumption data.
-Action Input: Share examples where changes in engine Displacement do not lead to proportional changes in Mpg.
-Observation: By examining specific fuel consumption data and case studies, it becomes evident that Displacement alone is not the sole determinant of Mpg, as other factors also contribute to fuel efficiency.
-Thought: I now know the final answer.
-Final answer: No
End Question number 4
Provide feedback to chat_manager. Press enter to skip and use auto-reply, or type exit to end the conversation: Judge, after receiving the responses from both sides, please make a final decision for each question
Admin (to chat_manager):
Judge, after receiving the responses from both sides, please make a final decision for each question
Judge (to chat_manager):
# Begin response of Judge #
Question number 1:
-Question: Are there any direct causal relationships from Displacement to Weight?
-Thought: Both the Affirmative_Causal_Agent and the Negative_Causal_Agent provided detailed explanations regarding the relationship between Displacement and Weight. They both highlighted that while engine Displacement can indirectly influence the Weight of a car, there is no direct causal relationship. The Affirmative_Causal_Agent supported their argument with empirical evidence and real-world examples, strengthening their position.
-Final Answer: No direct causal relationship from Displacement to Weight.
Question number 2:
-Question: Are there any direct causal relationships from Displacement to Horsepower?
-Thought: The Affirmative_Causal_Agent argued that there is a direct causal relationship between Displacement and Horsepower, supported by empirical data and specific engine specifications. In contrast, the Negative_Causal_Agent disagreed, stating that while there is a correlation, it is not a direct causal relationship. However, the Affirmative_Causal_Agents evidence and examples provided a more convincing argument.
-Final Answer: Yes, there is a direct causal relationship from Displacement to Horsepower.
Question number 3:
-Question: Are there any direct causal relationships from Displacement to Acceleration?
-Thought: The Affirmative_Causal_Agent claimed a direct causal relationship between Displacement and Acceleration, supported by the impact of engine size on power delivery. The Negative_Causal_Agent disagreed, stating that while Displacement can impact Acceleration, it is not a direct causal relationship. The Affirmative_Causal_Agents detailed explanation and real-world examples provided a stronger argument.
-Final Answer: Yes, there is a direct causal relationship from Displacement to Acceleration.
Question number 4:
-Question: Are there any direct causal relationships from Displacement to Mpg(miles per gallon)?
-Thought: Both sides agreed that there is no direct causal relationship between Displacement and Mpg. The Affirmative_Causal_Agent highlighted the influence of other factors such as engine technology and driving habits on fuel efficiency, while the Negative_Causal_Agent also supported this conclusion. Both sides provided examples and explanations that supported the lack of direct causality.
-Final Answer: No direct causal relationship from Displacement to Mpg(miles per gallon).
# End response of Judge #
B.0.2 Coding Agents Model
Phase 1: Plan Debating
Listing 7: Plan Debating responses
Admin (to chat_manager):
Task: find the causal relationship among variables in the dataset
Here are some information about the dataset: on automobile design and performance for cars in the 1970s and early 1980s, this dataset consists of the variables around the fuel consumption of cars. With five variables: Weight”, Displacement”, Horsepower”, Acceleration and Mpg”(miles per gallon)s
There is some sample of the data : Displacement Mpg Horsepower Weight Acceleration
0 307.0 18.0 130 3504 12.0
1 350.0 15.0 165 3693 11.5
2 318.0 18.0 150 3436 11.0
3 304.0 16.0 150 3433 12.0
4 302.0 17.0 140 3449 10.5
Dont need to understand the data, just give a plan of how to find the causal relationship among variables
The required output is a human-readable causal graph should be in a matrix for example with 3 variables:
[[0, 0, 1], [0, 0, 1 ], [1,0, 0]] and link format (Hot whether-> More ice scream sale, More people-> increase ice scream sale, increase ice scream sale ->Hot whether). If the size is 5 use 5x5 matrix and so on
Affirmative_Causal_Agent (to chat_manager):
-Question: How can we find the causal relationship among variables in the dataset?
-Thought: We need to choose an appropriate causal inference algorithm that can handle observational data and construct a causal graph representing the relationships among the variables.
-Action: Choose the appropriate causal inference algorithm for the dataset
-Action Input: Evaluate the pros and cons of each algorithm based on the dataset information
-Observation: Each algorithm has unique features that make it suitable for different types of data and research questions.
-Thought: Considering the dataset consists of observational data on automobile design and performance variables, we need an algorithm that can handle observational data and construct a causal graph efficiently.
-Action: Select the algorithm that best fits the requirements of the dataset
-Action Input: Analyze the pros and cons of each algorithm in the context of the dataset
-Observation: Each algorithm has advantages and limitations that need to be considered for causal inference in this specific dataset.
-Thought: Based on the dataset characteristics and the need to construct a causal graph among the variables, we should choose an algorithm that can handle observational data efficiently and provide a clear causal structure.
-Action: Propose the appropriate algorithm for causal inference in this dataset
-Action Input: Consider the strengths and limitations of each algorithm and select the one that aligns best with the dataset requirements
-Observation: The selected algorithm should be able to handle the variables related to automobile design and performance effectively to uncover the causal relationships.
-Thought: I now know the final answer
-Final Answer: The PC algorithm is the most appropriate choice for causal inference in this dataset as it can handle observational data efficiently and construct a causal network representing the relationships among the variables.
Negative_Causal_Agent (to chat_manager):
-Question: How can we find the causal relationship among variables in the dataset?
-Thought: We need to critically evaluate the proposed algorithm choice and provide a different perspective to ensure a comprehensive analysis.
-Action: Disagree with the previous answer and provide an alternative viewpoint.
-Action Input: Highlight the limitations of the PC algorithm and propose a different algorithm for causal inference in the dataset.
-Observation: By presenting a different algorithm choice, we can enrich the discussion and explore alternative approaches to causal inference.
-Thought: Its essential to consider the datasets characteristics and the specific requirements for causal inference to provide a well-rounded recommendation.
-Action: Analyze the pros and cons of another algorithm for causal inference.
-Action Input: Evaluate the strengths and limitations of the GES algorithm in the context of the dataset information.
-Observation: By considering a different algorithm, we can offer a diverse perspective on the most suitable approach for uncovering causal relationships in the dataset.
-Thought: After a thorough analysis of the dataset and the available causal inference algorithms, its crucial to present a well-justified recommendation.
-Action: Propose the alternative algorithm for causal inference in this dataset.
-Action Input: Highlight how the GES algorithm can address the specific characteristics of the dataset and provide valuable insights into causal relationships.
-Observation: By presenting a different algorithm choice, we can broaden the understanding of causal inference methods and their applicability to the given dataset.
-Thought: I now know the final answer
-Final Answer: The GES algorithm is a suitable choice for causal inference in this dataset as it offers efficient handling of large datasets and provides scalability, which is crucial for analyzing variables related to automobile design and performance.
Judge (to chat_manager):
-Question: What is the best approach to finding the causal relationship among variables in the dataset?
-Thought: Both debaters have presented different perspectives on the choice of causal inference algorithm. Its essential to evaluate the arguments and determine the most suitable approach for uncovering causal relationships in the given dataset.
-Action: Review the arguments presented by both debaters and assess the strengths and limitations of each proposed algorithm.
-Action Input: Examine the characteristics of the dataset and the requirements for causal inference.
-Observation: By comparing the pros and cons of the PC algorithm and the GES algorithm, we can determine which approach aligns best with the datasets needs.
-Thought: Its crucial to prioritize accuracy, efficiency, and scalability in selecting the causal inference algorithm for this dataset.
-Action: Evaluate the applicability of each algorithm in handling observational data related to automobile design and performance variables.
-Action Input: Consider how each algorithm constructs causal graphs and represents relationships among the variables.
-Observation: By analyzing the capabilities of the PC algorithm and the GES algorithm, we can determine which one is better suited for uncovering causal relationships in the dataset.
-Thought: After a comprehensive review of the arguments and algorithm characteristics, its time to make a decision based on the datasets requirements.
-Final Answer: The GES algorithm is the most suitable choice for causal inference in this dataset as it offers scalability and efficiency in handling variables related to automobile design and performance, providing valuable insights into causal relationships.
Provide feedback to chat_manager. Press enter to skip and use auto-reply, or type exit to end the conversation: Negative, please give step-by-step plan for the coder to implement, remember the output requirements, output in the format: step 1: ..., step 2:... , step N:...
Admin (to chat_manager):
Negative, please give step-by-step plan for the coder to implement, remember the output requirements, output in the format: step 1: ..., step 2:... , step N:...
Negative_Causal_Agent (to chat_manager):
-Question: What is the step-by-step plan for implementing the GES algorithm for causal inference in the dataset?
-Thought: To provide a detailed plan for the coder, we need to outline the specific steps involved in implementing the GES algorithm effectively.
-Action: Develop a comprehensive step-by-step plan for coding the GES algorithm.
-Action Input: Consider the requirements of the dataset and the algorithms characteristics.
-Step 1: Preprocessing
- Perform data preprocessing to handle missing values, normalize the data if necessary, and ensure data quality.
-Step 2: Implement GES Algorithm Initialization
- Set up the initial conditions for the GES algorithm, including defining the search space and initializing the score function.
-Step 3: Score Calculation
- Calculate the scores for potential causal relationships using the BIC score function to evaluate the goodness of fit and model complexity.
-Step 4: Greedy Equivalence Search
- Implement the Greedy Equivalence Search algorithm to explore the space of Markov equivalence classes and find the optimal causal structure.
-Step 5: Model Evaluation
- Evaluate the identified causal structure and assess its interpretability and relevance to the dataset.
-Step 6: Output Generation
- Generate the causal graph representing the relationships among the variables in a human-readable format as specified, linking variables based on the causal connections.
-Step 7: Interpretation
- Interpret the causal graph to understand the causal relationships among the variables and their implications for automobile design and performance.
-Step 8: Documentation
- Document the implementation steps, results, and any insights gained from the causal inference process for future reference.
-Step 9: Testing and Validation
- Test the implemented GES algorithm on the dataset, validate the results, and ensure the accuracy of the causal graph generated.
-Step 10: Optimization
- Optimize the algorithm implementation for efficiency and scalability, considering the size and complexity of the dataset.
-Thought: By providing a detailed step-by-step plan, the coder can effectively implement the GES algorithm for causal inference in the dataset.
-Final Answer: The step-by-step plan for coding the GES algorithm for causal inference in the dataset includes preprocessing, algorithm initialization, score calculation, Greedy Equivalence Search, model evaluation, output generation, interpretation, documentation, testing and validation, and optimization steps to ensure accurate and insightful causal relationships among the variables related to automobile design and performance.
Phase 2: Code Executing
Listing 8: Code executing responses
ask: find the causal relationship among variables in the dataset
Here is the datasets directory: benchmark datasets/DWD climate data.csv
Here are some information about the dataset: on global climate variables , This dataset encompasses six continuous variables capturing climate observations such as altitude, temperature, precipitation levels, longitude, sunshine duration, and latitude. It is aimed at studying weather patterns, climate change impacts, and geographical correlations in climate variables.
Here is step-by-step plan for implement the code to find the causal relationship among variables:
1. **Data Preparation**:
- Load the dataset containing variables: Weight, Displacement, Horsepower, Acceleration, Mpg.
- Handle missing values and ensure data cleanliness.
2. **Graph Initialization**:
- Create an initial fully connected, undirected graph with all variables connected to each other.
3. **Conditional Independence Testing**:
- Implement statistical tests (e.g., chi-squared tests, correlation tests) to check for conditional independence between variable pairs given a set of other variables.
- Remove edges where conditional independence is observed.
4. **Orientation Rules**:
- Apply orientation rules to determine the directionality of edges based on conditional independence patterns.
- Convert the undirected graph into a directed acyclic graph (DAG).
5. **Iteration**:
- Iterate through increasing sizes of conditioning sets (empty set, singletons, pairs, etc.) to refine the graph structure.
- Remove edges based on conditional independence tests until no more edges can be deleted.
6. **Implementation Steps**:
- Develop functions or modules for graph construction, conditional independence testing, orientation rules, and iterative edge removal.
- Utilize appropriate statistical libraries for conducting tests and graph manipulation.
7. **Execution and Evaluation**:
- Run the implemented PC algorithm on the preprocessed dataset.
- Evaluate the resulting causal graph for meaningful relationships and causal directions.
- Validate the graph structure and assess the strength of causal connections.
8. **Documentation and Reporting**:
- Document the code implementation, including functions, algorithms used, and any modifications made.
- Generate a human-readable causal graph matrix based on the algorithms output.
- Prepare a detailed report summarizing the causal relationships discovered and their implications.
The required output is a human-readable causal graph should be in a matrix for example with 3 variables:
[[0, 0, 1], [0, 0, 1 ], [1,0, 0]] and link format (Hot whether-> More ice scream sale, More people-> increase ice scream sale, increase ice scream sale ->Hot whether). If the size is 5 use 5x5 matrix and so on
Code_Writer (to chat_manager):
-Question: What is the initial step to find the causal relationship among variables in the dataset?
-Thought: The initial step is to load the dataset and handle missing values to ensure data cleanliness.
-Action: Load the dataset and handle missing values
-Action Input:
‘‘‘python
import pandas as pd
# Load the dataset
data = pd.read_csv("benchmark datasets/DWD climate data.csv")
# Check for missing values
missing_values = data.isnull().sum()
missing_values
-Observation: The dataset has been loaded and missing values have been checked.
-Question: What is the next step after loading the dataset and handling missing values?
-Thought: The next step is to create an initial fully connected, undirected graph with all variables connected to each other.
-Action: Create an initial fully connected, undirected graph.
-Action Input:
‘‘‘python
# Create an initial fully connected, undirected graph
import numpy as np
num_variables = len(data.columns)
initial_graph = np.ones((num_variables, num_variables))
initial_graph
-Observation: The initial fully connected, undirected graph has been created.
-Question: What is the next step after creating the initial fully connected, undirected graph?
-Thought: The next step is to implement statistical tests to check for conditional independence between variable pairs given a set of other variables.
-Action: Implement statistical tests for conditional independence.
-Action Input: Implementing statistical tests for conditional independence.
-Question: What is the next step after implementing statistical tests for conditional independence?
-Thought: The next step is to apply orientation rules to determine the directionality of edges based on conditional independence patterns and convert the graph into a directed acyclic graph (DAG).
-Action: Apply orientation rules and convert the graph into a DAG.
-Action Input: Applying orientation rules based on conditional independence patterns.
We will continue with the rest of the steps after applying orientation rules and converting the graph into a DAG.
>>>>>>>> EXECUTING 2 CODE BLOCKS (inferred languages are [python, python])...
Code_Executor (to chat_manager):
exitcode: 0 (execution succeeded)
Code output: Altitude 0
Temperature 0
Precipitation 0
Longitude 0
Sunshine 0
Latitude 0
dtype: int64
array([[1., 1., 1., 1., 1., 1.],
[1., 1., 1., 1., 1., 1.],
[1., 1., 1., 1., 1., 1.],
[1., 1., 1., 1., 1., 1.],
[1., 1., 1., 1., 1., 1.],
[1., 1., 1., 1., 1., 1.]])
Code_Writer (to chat_manager):
-Question: What is the next step after implementing statistical tests for conditional independence and handling missing values?
-Thought: The next step is to apply orientation rules to determine the directionality of edges based on conditional independence patterns and convert the graph into a directed acyclic graph (DAG).
-Action: Apply orientation rules and convert the graph into a DAG.
-Action Input: Applying orientation rules based on conditional independence patterns.
‘‘‘python
# We will continue with this step after applying orientation rules and converting the graph into a DAG
# The next step is to implement the PC Algorithm for causal relationship discovery.
>>>>>>>> EXECUTING CODE BLOCK (inferred language is python)...
Code_Executor (to chat_manager):
exitcode: 0 (execution succeeded)
Code output:
Code_Writer (to chat_manager):
-Question: What is the next step after applying orientation rules and converting the graph into a DAG?
-Thought: The next step is to implement the PC Algorithm for causal relationship discovery.
-Action: Implement the PC Algorithm for causal relationship discovery.
-Action Input: Implementing the PC Algorithm with the custom parameters if needed.
Code_Executor (to chat_manager):
Code_Writer (to chat_manager):
-Question: What is the next step after implementing the PC Algorithm for causal relationship discovery?
-Thought: The next step is to create a human-readable causal graph matrix based on the algorithms output.
-Action: Create a human-readable causal graph matrix.
-Action Input:
‘‘‘python
# Using the provided create_adjacency_matrix function to get the causal graph matrix
def create_adjacency_matrix(cg):
num_nodes = len(cg.G.nodes)
adj_matrix = np.zeros((num_nodes, num_nodes), dtype=int)
for i in range(num_nodes):
for j in range(num_nodes):
# i <- j
if cg.G.graph[i][j] == 1 and cg.G.graph[j][i] == -1:
adj_matrix[i, j] = 1
# i -- j
elif cg.G.graph[i][j] == -1 and cg.G.graph[j][i] == -1:
adj_matrix[i, j] = -1
# i <->
elif cg.G.graph[i][j] == 1 and cg.G.graph[j][i] == 1:
adj_matrix[i, j] = 2
return adj_matrix
# Get adjacency matrix from the PC Algorithm output
total_adj_matrix_pc = create_adjacency_matrix(cg)
total_adj_matrix_pc
This will provide a human-readable causal graph matrix based on the output of the PC Algorithm.
>>>>>>>> EXECUTING CODE BLOCK (inferred language is python)...
Code_Executor (to chat_manager):
exitcode: 1 (execution failed)
Code output: ERROR: ERROR: NameError: name cg is not defined
[…]
Code_Writer (to chat_manager):
-Question: What can be done to resolve the NameError and successfully create a human-readable causal graph matrix?
-Thought: We need to ensure that the variable cg‘, which holds the output of the PC Algorithm, is properly defined before calling the function to create the adjacency matrix.
-Action: Execute the PC Algorithm to obtain the graph object cg before creating the adjacency matrix.
-Action Input:
‘‘‘python
from causallearn.search.ConstraintBased.PC import pc
# Assuming data is prepared and available
# Execute the PC Algorithm to obtain the causal graph cg
cg = pc(data)
# Function to create the adjacency matrix
def create_adjacency_matrix(cg):
num_nodes = len(cg.G.nodes)
adj_matrix = np.zeros((num_nodes, num_nodes), dtype=int)
for i in range(num_nodes):
for j in range(num_nodes):
# i <- j
if cg.G.graph[i][j] == 1 and cg.G.graph[j][i] == -1:
adj_matrix[i, j] = 1
# i -- j
elif cg.G.graph[i][j] == -1 and cg.G.graph[j][i] == -1:
adj_matrix[i, j] = -1
# i <->
elif cg.G.graph[i][j] == 1 and cg.G.graph[j][i] == 1:
adj_matrix[i, j] = 2
return adj_matrix
# Obtain the adjacency matrix from the PC Algorithm output
total_adj_matrix_pc = create_adjacency_matrix(cg)
total_adj_matrix_pc
>>>>>>>> EXECUTING CODE BLOCK (inferred language is python)...
Code_Executor (to chat_manager):
exitcode: 1 (execution failed)
Code output: ERROR: ERROR: AssertionError: Input data must be a numpy array.
[…]
Code_Writer (to chat_manager):
-Question: How can we resolve the Assertion Error regarding the input data format in order to successfully create the adjacency matrix for the causal graph?
-Thought: The error indicates that the input data must be in the format of a NumPy array. We need to convert the data into a NumPy array before using it in the PC Algorithm.
-Action: Convert the dataset to a NumPy array before executing the PC Algorithm.
-Action Input:
‘‘‘python
import numpy as np
# Convert the Pandas DataFrame to a NumPy array
data_np = data.to_numpy()
# Execute the PC Algorithm with the converted NumPy array data
cg = pc(data_np)
# Function to create the adjacency matrix
def create_adjacency_matrix(cg):
num_nodes = len(cg.G.nodes)
adj_matrix = np.zeros((num_nodes, num_nodes), dtype=int)
for i in range(num_nodes):
for j in range(num_nodes):
# i <- j
if cg.G.graph[i][j] == 1 and cg.G.graph[j][i] == -1:
adj_matrix[i, j] = 1
# i -- j
elif cg.G.graph[i][j] == -1 and cg.G.graph[j][i] == -1:
adj_matrix[i, j] = -1
# i <->
elif cg.G.graph[i][j] == 1 and cg.G.graph[j][i] == 1:
adj_matrix[i, j] = 2
return adj_matrix
# Obtain the adjacency matrix from the PC Algorithm output
total_adj_matrix_pc = create_adjacency_matrix(cg)
total_adj_matrix_pc
By converting the dataset to a NumPy array before executing the PC Algorithm, the Assertion Error regarding the input data format can be resolved.
>>>>>>>> EXECUTING CODE BLOCK (inferred language is python)...
Code_Executor (to chat_manager):
exitcode: 0 (execution succeeded)
Code output: 0%| | 0/6 [00:00<?, ?it/s]
{"version_major": 2, "version_minor": 0, "model_id": "a1056c8cce5046d49d2ba46d506c5bc8"}
array([[ 0, 1, 0, 0, 0, 1],
[ 0, 0, 0, 0, 0, 0],
[ 1, 0, 0, 0, 0, 0],
[ 0, 1, 0, 0, 1, 0],
[ 0, 0, 0, 0, 0, -1],
[ 0, 0, 0, 0, -1, 0]])
Code_Writer (to chat_manager):
-Question: What is the adjacency matrix for the human-readable causal graph based on the output of the PC Algorithm?
-Observation: The adjacency matrix for the human-readable causal graph based on the output of the PC Algorithm is as follows:
[[ 0, 1, 0, 0, 0, 1],
[ 0, 0, 0, 0, 0, 0],
[ 1, 0, 0, 0, 0, 0],
[ 0, 1, 0, 0, 1, 0],
[ 0, 0, 0, 0, 0, -1],
[ 0, 0, 0, 0, -1, 0]]
This adjacency matrix represents the causal relationships among the variables: Altitude, Temperature, Precipitation, Longitude, Sunshine, and Latitude. The numbers in the matrix indicate the directionality of the edges between the variables.
>>>>>>>> EXECUTING CODE BLOCK (inferred language is python)...
Code_Executor (to chat_manager):
exitcode: 0 (execution succeeded)
Code output: [[0, 1, 0, 0, 0, 1],
[0, 0, 0, 0, 0, 0],
[1, 0, 0, 0, 0, 0],
[0, 1, 0, 0, 1, 0],
[0, 0, 0, 0, 0, -1],
[0, 0, 0, 0, -1, 0]]
Code_Writer (to chat_manager):
-Question: What is the adjacency matrix for the human-readable causal graph based on the output of the PC Algorithm?
-Final Answer: The adjacency matrix for the human-readable causal graph based on the output of the PC Algorithm is:
[[0, 1, 0, 0, 0, 1],
[0, 0, 0, 0, 0, 0],
[1, 0, 0, 0, 0, 0],
[0, 1, 0, 0, 1, 0],
[0, 0, 0, 0, 0, -1],
[0, 0, 0, 0, -1, 0]]
This matrix represents the causal relationships among the variables: Altitude, Temperature, Precipitation, Longitude, Sunshine, and Latitude.
>>>>>>>> EXECUTING CODE BLOCK (inferred language is python)...
Code_Executor (to chat_manager):
exitcode: 0 (execution succeeded)
Code output: [[0, 1, 0, 0, 0, 1],
[0, 0, 0, 0, 0, 0],
[1, 0, 0, 0, 0, 0],
[0, 1, 0, 0, 1, 0],
[0, 0, 0, 0, 0, -1],
[0, 0, 0, 0, -1, 0]]
Code_Writer (to chat_manager):
Final Answer: Based on the output of the PC Algorithm on the provided climate dataset, the human-readable causal graph matrix is as follows:
[[0, 1, 0, 0, 0, 1],
[0, 0, 0, 0, 0, 0],
[1, 0, 0, 0, 0, 0],
[0, 1, 0, 0, 1, 0],
[0, 0, 0, 0, 0, -1],
[0, 0, 0, 0, -1, 0]]
This matrix indicates the causal relationships among the variables in the dataset: Altitude, Temperature, Precipitation, Longitude, Sunshine, and Latitude.
>>>>>>>> EXECUTING CODE BLOCK (inferred language is python)...
Code_Executor (to chat_manager):
exitcode: 0 (execution succeeded)
Code output: [[0, 1, 0, 0, 0, 1],
[0, 0, 0, 0, 0, 0],
[1, 0, 0, 0, 0, 0],
[0, 1, 0, 0, 1, 0],
[0, 0, 0, 0, 0, -1],
[0, 0, 0, 0, -1, 0]]
Code_Writer (to chat_manager):
Final Results:
- The adjacency matrix for the human-readable causal graph based on the output of the PC Algorithm is:
[[0, 1, 0, 0, 0, 1],
[0, 0, 0, 0, 0, 0],
[1, 0, 0, 0, 0, 0],
[0, 1, 0, 0, 1, 0],
[0, 0, 0, 0, 0, -1],
[0, 0, 0, 0, -1, 0]]
This matrix represents the causal relationships among the variables Altitude, Temperature, Precipitation, Longitude, Sunshine, and Latitude in the climate dataset. The numbers indicate the directionality of the causal edges in the graph.
FINISH
Admin (to chat_manager):
export the matrix into csv file call "Auto_Coding_MAD_React.csv"
Code_Writer (to chat_manager):
‘‘‘python
import pandas as pd
# Create the adjacency matrix as a Pandas DataFrame
adj_matrix_data = [[0, 1, 0, 0, 0, 1],
[0, 0, 0, 0, 0, 0],
[1, 0, 0, 0, 0, 0],
[0, 1, 0, 0, 1, 0],
[0, 0, 0, 0, 0, -1],
[0, 0, 0, 0, -1, 0]]
columns = [’Altitude’, Temperature’, Precipitation’, Longitude’, Sunshine’, Latitude’]
adj_matrix_df = pd.DataFrame(adj_matrix_data, columns=columns, index=columns)
# Export the adjacency matrix to a CSV file
adj_matrix_df.to_csv("Auto_Coding_MAD_React.csv")
>>>>>>>> EXECUTING CODE BLOCK (inferred language is python)...
Code_Executor (to chat_manager):
exitcode: 0 (execution succeeded)
Code output:
Code_Writer (to chat_manager):
The adjacency matrix representing the causal relationships among the variables Altitude, Temperature, Precipitation, Longitude, Sunshine, and Latitude has been successfully exported to a CSV file named "Auto_Coding_MAD_React.csv".
Code_Executor (to chat_manager):
Code_Writer (to chat_manager):
I have successfully completed all the steps to find the causal relationship among variables in the climate dataset:
1. Data Preparation: Loaded the dataset and handled missing values.
2. Graph Initialization: Created an initial fully connected, undirected graph.
3. Conditional Independence Testing: Implemented statistical tests to check for conditional independence between variable pairs.
4. Orientation Rules: Applied orientation rules to determine the directionality of edges based on conditional independence patterns and converted the graph into a directed acyclic graph (DAG).
5. Implemented the PC Algorithm for causal relationship discovery and obtained the adjacency matrix for the human-readable causal graph.
The adjacency matrix has been exported to a CSV file named "Auto_Coding_MAD_React.csv".
I have provided a comprehensive solution to find the causal relationships in the climate dataset. If you have any more tasks or questions, feel free to let me know!
Code_Executor (to chat_manager):
Code_Writer (to chat_manager):
FINISH